### v4

# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

Abstract
Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL). Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value. Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces. To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs). Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation. A custom-designed coder-evaluator framework is further employed to guide the generation and verification of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference. Furthermore, an IGM-Gating Mechanism enables QLLM to flexibly enforce or relax the monotonicity constraint depending on task demands, covering both IGM-compliant and non-monotonic scenarios. Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios. The code is available at https://github.com/zhouyangjiang71-sys/QLLM.
Introduction
Recent advances in cooperative multi-agent reinforcement learning (MARL)(Oroojlooy and Hajinezhad 2023) have led to widespread applications in domains such as assembly lines (Kaven et al. 2024), logistics distribution (Yan et al. 2022), autonomous driving (Zhang et al. 2024b), and swarm robotics (Orr and Dutta 2023). In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training. A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called ”lazy agents” (Liu et al. 2023).
To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community. Attention-based credit assignment methods (Soydaner 2022) incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions. By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions. This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
Figure 1: A comparison between the traditional value decomposition method using a mixing network (Left) and our proposed novel paradigm QLLM leveraging LLMs (Right). The traditional approach employs a neural network to model the nonlinear relationship between local Q-values and the global Q-value, whereas QLLM capitalizes on the extensive knowledge encoded within LLMs to directly generate a training-free credit assignment function.
Policy gradient approaches (Agarwal et al. 2021) tackle credit assignment from a differentiable optimization perspective. Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability. These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation. In addition, value decomposition methods (Sunehag et al. 2017) have become increasingly popular. These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent. The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution. These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for each agents. A representative example of this is the mixing network employed in value decomposition methods. These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability (Linardatos, Papastefanopoulos, and Kotsiantis 2020). With the advancement of large language models (LLMs) (Kasneci et al. 2023), new paradigms have emerged to address many long-standing challenges in MARL. By leveraging extensive pre-trained knowledge (Petroni et al. 2019), superior generalization capacity (Anil et al. 2022), and powerful code generation abilities (Xu et al. 2022a), LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines. Recent studies (Qu et al. 2025) have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals. Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines. Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency (Sanh et al. 2021). In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality (Zhang et al. 2023b). Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network. To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation. Within this framework, LLMs act as both a code generator and a code evaluator (Desmond et al. 2024). Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT (Rashid et al. 2020b; Su, Adams, and Beling 2021; Hu et al. 2021). Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability. Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues (Martino, Iannelli, and Truong 2023) commonly encountered by LLMs. This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network. Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks. An IGM-Gating Mechanism is further introduced to flexibly switch the monotonicity constraint on or off, adapting to tasks where the Individual-Global-Max principle holds or does not. The main contributions of this study are summarized as follows:
  1. 1.
    We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF. This framework is generalizable and can be extended to other MARL scenarios enhanced by LLMs.
  2. 2.
    QLLM is proposed by combining TFCAF and value function decomposition while remaining compatible with most MARL algorithms that employ mixing networks for credit assignment. Unlike existing methods, QLLM can accurately infer each agent’s contribution without interacting with the environment.
  3. 3.
    Extensive experiments demonstrate that QLLM outperforms existing credit assignment algorithms, particularly in scenarios involving high-dimensional state spaces. Moreover, the integration of LLMs endows the contribution assignment among agents with semantic interpretability.
Related Work
Credit Assignment in MARL
In recent years, the credit assignment problem has emerged as a fundamental challenge in multi-agent systems. And policy-based approaches (Yu et al. 2022; Zhang et al. 2024a) were among the first to introduce credit assignment mechanisms. MADDPG (Lowe et al. 2017) extends the DDPG (Lillicrap et al. 2015) framework to multi-agent settings by incorporating centralized training with decentralized execution. This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments. MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA (Foerster et al. 2018) leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG. MAAC (Iqbal and Sha 2019) further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
Another widely adopted paradigm is the value decomposition approach (Rashid et al. 2020b; Xu et al. 2022b), which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN (Sunehag et al. 2017) models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX (Rashid et al. 2020b) improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN (Son et al. 2019) and QPLEX (Wang et al. 2020) mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. LICA (Zhou et al. 2020) adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information. Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks. RIIT (Hu et al. 2021) goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training. Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
LLM-enhanced Reinforcement Learning
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning (Zhang et al. 2023a; Li et al. 2024). In the SMAC-R1 framework (Deng et al. 2024), agents utilize LLMs to generate decision tree code by providing task descriptions. The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards. Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments. They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes. For instance, the Reflexion framework (Shinn et al. 2023) introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters. Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions. Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks. The EUREKA framework (Ma et al. 2023) achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
These examples demonstrate the powerful information processing and code generation capabilities of LLMs. Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL. Qu et al. (2025) were the first to introduce LLMs into the credit assignment problem. They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly. This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms. However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components. Additionally, Lin et al. (2025) employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions. These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance. Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
Preliminaries
Decentralized Partially Observable Markov Decision Process
In this work, we focus on a fully cooperative MARL setting that can be formalized as a decentralized partially observable Markov decision process (Dec-POMDP). A Dec-POMDP is typically defined by a tuple $G=\langle S,A,P,r,Z,O,n,\gamma\rangle$. Here, $n$ agents indexed by $i\in\{1,\dots,n\}$ operate in an environment with true underlying state $s\in S$. At each timestep, agent $i$ selects an action $a^{i}\in A$ based solely on its private observation $z_{i}\in Z$, which is sampled according to the observation function $O(s,i)$. The joint action of all agents is denoted by $\boldsymbol{a}\in\boldsymbol{A}$.
The environment evolves according to the transition probability function $P(s^{\prime}\mid s,\boldsymbol{a}):S\times\boldsymbol{A}\to[0,1]$, which specifies the likelihood of transitioning to state $s^{\prime}$ given the current state $s$ and joint action $\boldsymbol{a}$. All agents cooperate to maximize a shared reward signal defined by the function $r(s,\boldsymbol{a}):S\times\boldsymbol{A}\to\mathbb{R}$. The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward $\sum_{j=0}^{\infty}\gamma^{j}r_{t+j}$, where $\gamma\in[0,1)$ is a discount factor determining the weight of future rewards.
A key challenge in Dec-POMDPs lies in the credit assignment problem. This issue arises because all agents within the system share a common reward structure, making it inherently difficult to accurately attribute individual contributions to the collective outcome. Furthermore, the environment is perceived as non-stationary from the viewpoint of each agent. This non-stationarity poses significant challenges in achieving effective coordination and optimal decision-making, as agents must adapt to the changing strategies of others while managing incomplete state representations.
Figure 2: QLLM is designed within a coder-evaluator framework to autonomously generate high-quality, training-free credit assignment functions (TFCAFs). In each iteration, the coder LLM $M_{\text{coder}}$ produces $K$ candidate functions based on a task prompt and a role-specific instruction, referred to as the coder prompt. These candidate functions are initially assessed using environmental state information and local Q-values. If any runtime or syntax errors are identified, the faulty candidates are discarded and regenerated. Subsequently, the evaluator LLM $M_{\text{evaluator}}$ reviews the candidate functions, selects the most promising one based on its evaluator prompt This process is generated by LLMs through structured prompts without any need for human intervention.
Value Decomposition and IGM Principle
Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions $Q_{i}$ of individual agents into a global Q-value function $Q_{\text{tot}}$. The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint. The IGM principle (Hong, Jin, and Tang 2022) requires that the joint action that maximizes the global action-value function $Q_{\text{tot}}$ must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
$\arg\max_{\boldsymbol{a}}Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{a})=\left(\begin{array}[]{c}\arg\max_{a^{1}}Q_{1}\left(\tau^{1},a^{1}\right)\\ \vdots\\ \arg\max_{a^{n}}Q_{n}\left(\tau^{n},a^{n}\right)\end{array}\right),$ (1)
where $\boldsymbol{\tau}\in T^{n}$ represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this, Qatten (Yang et al. 2020) provides a theoretical extension, showing that $Q_{tot}$ can be expressed as a linear combination of the individual $Q_{i}$ functions:
$Q_{\text{tot}}(s,\boldsymbol{a})=\sum_{i=1}^{n}\sum_{h=1}^{H}w_{i,h}(s)Q_{i}(\tau^{i},a^{i})+b.$ (2)
Here, $w_{i,h}(s)$ denotes a coefficient that depends on the state $s$, $H$ is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of $Q_{\text{tot}}$ with respect to $Q_{i}$, where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on $Q_{\text{tot}}$ is sufficiently small to be ignored or effectively absorbed into a bias term $b$, allowing for a simplified model formulation and reduced computational complexity. Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
Method
This section presents a detailed discussion of the implementation of each component of QLLM. We first propose the concept of the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and more relaxed joint value modeling. This provides a foundation for adaptively handling different types of multi-agent tasks. Then we introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination (Tonmoy et al. 2024) and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
IGM-Gating Mechanism
To flexibly handle diverse coordination demands across tasks, we introduce the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and relaxed joint value modeling. This mechanism allows for conditional enforcement of the monotonicity constraint in the credit assignment.
Definition 1 (IGM-Gating Mechanism).
Let $Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})$ be the joint action-value function, and $\{Q_{i}(s_{i},a_{i})\}_{i=1}^{n}$ the individual utility functions. Define a binary gating variable $\gamma\in\{0,1\}$, where $\gamma=1$ indicates that IGM is enforced. The IGM-Gating Mechanism imposes the following conditional monotonicity constraint:
$\forall i,\quad\gamma=1\Longrightarrow\frac{\partial Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})}{\partial Q_{i}(s_{i},a_{i})}\geq 0.$ (3)
When the gating variable is inactive ($\gamma=0$), no monotonicity constraint is enforced, allowing $Q_{\text{tot}}$ to freely capture non-monotonic dependencies.
Training-Free Credit Assignment Function
As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions. The network takes as input the individual Q-values of all agents, denoted as $\boldsymbol{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$, and produces the joint Q-value $Q_{\text{tot}}$ as output. To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism (Ha, Dai, and Le 2017). Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state $s$. This design facilitates state-aware adaptation while ensuring compatibility with CTDE framework, making deployment straightforward. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
To address these issues, we propose the Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM. This module is generated by LLMs through iterative error correction, based on a series of easily accessible task-related prompts. As illustrated in Equation (2), the global Q-value function $Q_{\text{tot}}$ can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF, which we formally define below.
Definition 2.
A Training-Free Credit Assignment Function (TFCAF) is defined as:
$Q_{\text{tot}}\left(s,\boldsymbol{a}\right)=\sum_{i=1}^{n}f_{w}^{i}(s)\,Q_{i}\left(\tau^{i},a^{i}\right)+f_{b}(s),$ (4)
where $f_{w}^{i}(s)$ and $f_{b}(s)$ are general nonlinear functions of the global state $s$ generated by LLMs before training.
As illustrated in Definition 1, when the gating signal $\gamma=1$, the function $f_{w}^{i}(s)$ is strictly constrained to be non-negative. In contrast, when $\gamma=0$, this constraint is lifted. Benefiting from the rich prior knowledge of LLMs, the appropriate value of $\gamma$ can be easily inferred from the task description, enabling the model to adapt to different types of tasks. For instance, in matrix game scenarios, where individual optimal decisions do not necessarily align with the global optimum, the LLM may assign a negative value to $f_{w}^{i}(s)$. The coefficients $f_{w}^{i}\left(s\right)$ and the bias term $f_{b}\left(s\right)$ make the overall TFCAF a nonlinear mapping, although its final combination with the local $Q_{i}$ remains linear in form. Both components are automatically produced by LLMs through the coder-evaluator framework, which we detail next along with the iterative correction and its role in credit assignment for multi-agent systems.
Coder Evaluator Framework
Due to the susceptibility of LLMs to hallucinations, the process of generating TFCAF must be designed to be robust. To address this challenge, we propose the coder-evaluator framework, which automatically generates TFCAF through a modular and self-correcting process, ensuring both accurate credit assignment and logical consistency in code generation.
This framework employs two LLMs with distinct yet complementary roles: the coder LLM $M_{coder}$, responsible for generating candidate TFCAFs, and the evaluator LLM $M_{evaluator}$, which critically evaluates these candidates. All of these are generated by LLMs through structured prompts without any need for human intervention.
The interaction between the two models is driven by two types of prompts: task prompts ($P_{task}$) specify the meanings of each dimension in the global state space, the action space of the agents, and the composition of the reward function. This information is objective and can be easily obtained from the environment; role prompts ($P_{role}$) define the fixed responsibilities and operational roles of each LLM within the framework. The role prompts are further categorized into coder prompts and evaluator prompts, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively. This clear separation and stability of prompt design make QLLM easy to adapt to new cooperative tasks without extensive manual prompt tuning or subjective rewriting (see in Appendix A).
The generation of TFCAF can be formally described through the following three clearly-defined steps.
Initial Generation.
Firstly, $M_{coder}$ generates $K$ candidate credit assignment functions by processing both the role-specific and task-specific prompts:
$\phi_{1},\phi_{2},...,\phi_{K}\ \Leftarrow\ M_{coder}\big(P_{role},\ P_{task}\big),$ (5)
where $\phi_{1},\phi_{2},\dots,\phi_{K}$ denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases.
Error Detection and Correction.
Due to the inherent hallucination tendencies of LLMs, candidate functions might contain syntax or dimensional errors. To address this, each candidate TFCAF is compiled and run once with any available input state and local Q-values. If an error occurs during execution, the error message generated by the compiler will be fed back to $M_{coder}$, prompting regeneration of the faulty TFCAF:
$\phi_{correct}\,\,\Leftarrow\,\,M_{coder}\big(P_{role},\,\,P_{task},\ error\big).$ (6)
This correction step repeats until all $K$ candidate TFCAFs can execute successfully without errors.
Selection of TFCAF.
Subsequently, $M_{evaluator}$ evaluates these validated candidate TFCAFs, selecting the most promising one based on task relevance, interpretability, and code clarity. Notably, the selection of the best candidate is fully guided by structured prompts and performed automatically by $M_{evaluator}$, without any manual or subjective intervention:
$choice\,\Leftarrow\,\,M_{evaluator}\big(P_{role},\ P_{task},\ {\Phi}\big),$ (7)
where ${\Phi}=\{\phi_{1},\phi_{2},...,\phi_{K}\}$.
Algorithm 1 Pseudo-code for QLLM
1:Part 1: TFCAF Generation (LLM-based)
2: $\phi_{1},\phi_{2},\dots,\phi_{K}\leftarrow M_{\text{coder}}(P_{\text{role}},\,P_{\text{task}})$ $\triangleright$ Generate $K$ candidate functions
3:for all $\phi_{k}\in\{\phi_{1},\dots,\phi_{K}\}$ do
4:  if $\phi_{k}$ fails on environment samples then
5:   $\phi_{k}\leftarrow M_{\text{coder}}(P_{\text{role}},P_{\text{task}},\text{error})$ $\triangleright$ Regenerate function based on error message
6:  end if
7:end for
[…]
10:Generate $f_{\text{TFCAF}}$ from $\phi_{\text{final}}$ using predefined template
11:Part 2: Agents Optimization (RL-based)
12:for episode = 1 to $M$ do
13:  for each agent $i$ do
14:   $a^{i}\leftarrow\arg\max_{a}Q_{i}(\tau^{i},a;\theta)$
15:  end for
16:  Compute global Q-value:
17:   $Q_{\text{tot}}(s,\boldsymbol{a})=f_{\text{TFCAF}}(Q_{1},...,Q_{n};s)$
18:  Store transition $(s,\boldsymbol{a},r,s^{\prime})$ in buffer $\mathcal{D}$
19:  Sample batch and compute TD target:
20:   $y=r+\gamma\max_{\boldsymbol{a}^{\prime}}Q_{\text{tot}}(s^{\prime},\boldsymbol{a}^{\prime};\theta^{-})$
21:  Update parameters $\theta$ by minimizing the TD loss:
22:   $\mathcal{L}(\theta)=\mathbb{E}\left[\left(y-Q_{\text{tot}}(s,\boldsymbol{a};\theta)\right)^{2}\right].$
23:end for
Training Procedure
With the coder-evaluator framework, we obtain a training-free credit assignment function. Like QMIX and VDN, QLLM is a value-based MARL method developed under the CTDE paradigm. Its objective is to learn a global action-value function $Q_{\text{tot}}(s,\boldsymbol{a})$ that can be decomposed into individual utility functions $Q_{i}(\tau^{i},a^{i})$, while avoiding the constraints imposed by conventional hypernetwork-based mixing networks. This design enables the LLM to model more complex and nonlinear interactions between local and global Q-values, allowing it to better adapt to the specific structural properties of the task.
As with most value decomposition approaches, each agent $i$ in QLLM employs a deep neural network to approximate its individual action-value function $Q_{i}(\tau^{i},a^{i};\theta)$, where $\theta$ denotes the set of learnable parameters. These individual Q-values are then aggregated by the TFCAF, which is automatically generated by LLMs, to produce the global action-value function:
$Q_{\text{tot}}(s,\boldsymbol{a})=f_{\text{TFCAF}}(Q_{1},\dots,Q_{n};s),$ (8)
where $f_{\text{TFCAF}}$ is a parameterized function in which the weights are allowed to take arbitrary values, while the biases are also unconstrained. $f_{\text{TFCAF}}$ receives the agent-specific Q-values $Q_{i}$ as input and utilizes the global state $s$ to generate the corresponding weights and biases through the weight module $f_{w}^{i}$ and the bias module $f_{b}$, respectively. Notably, all parameters in the TFCAF are generated by LLMs and remain fixed during training. The global Q-value is then computed as shown in Equation (4). The training objective is to minimize the temporal-difference (TD) loss between the predicted total Q-value and the target Q-value:
$\mathcal{L}(\theta)=\mathbb{E}_{(s,\boldsymbol{a},r,s^{\prime})}\bigg[\big(y-Q_{\text{tot}}(s,\boldsymbol{a};\theta)\big)^{2}\bigg],$ (9)
where the target value $y$ is computed using Bellman equation:
$y=r+\gamma\max_{\boldsymbol{a}^{\prime}}Q_{\text{tot}}(s^{\prime},\boldsymbol{a}^{\prime};\theta^{-}),$ (10)
with $\gamma$ being the discount factor. $\theta^{-}$ denotes the parameters of the target agent network.
Compared to other neural network-based approaches for multi-agent credit assignment, QLLM involves fewer trainable parameters. This facilitates more precise credit allocation among agents in the early training stages, thereby accelerating the convergence of their decision-making policies. Meanwhile, the incorporation of TFCAF allows QLLM to effectively exploit prior knowledge while employing an interpretable credit assignment mechanism.
Experiments
Figure 3: Average episodic return curves for selected tasks in GRF and MPE environments (the results for LBF are shown in Appendix C).
To comprehensively evaluate the proposed method, the experiments were designed with six main objectives: (1) to assess the performance of the QLLM algorithm in terms of average return across multi-agent environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other mixing-based algorithms; (4) to demonstrate the compatibility of QLLM with various LLMs; (5) to validate the effectiveness of the proposed coder-evaluator framework under various parameters; (6) to demonstrate the superiority of our method in terms of training costs.
Experimental Setups
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary LLM used in our method is the DeepSeek-R1 (Guo et al. 2025) inference model. For each experiment, the reported return values represent the mean over five independent training runs, with 90% confidence intervals indicated by error bars. Unless otherwise specified, all environment parameters are set to their default values. Our experimental environments include: (1) Cooperative Matrix Games [^1] https://github.com/uoe-agents/matrix-games, used to validate the algorithm’s performance in non-monotonic scenarios, which may violate the IGM principle; (2) Multi-agent Particle Environments (MPE) [^2] https://pettingzoo.farama.org/environments/mpe/, for evaluating performance in high-dimensional state spaces; (3) Level-Based Foraging (LBF) [^3] https://github.com/uoe-agents/lb-foraging, for testing in sparse reward settings; and (4) Google Research Football (GRF) [^4] https://github.com/google-research/football, to assess performance in complex, realistic environments.
We compare QLLM with six commonly used baseline algorithms: QMIX (Rashid et al. 2020b), QPLEX (Wang et al. 2020), Weighted QMIX (Rashid et al. 2020a), Qatten (Yang et al. 2020), RIIT (Hu et al. 2021), MASER (Jeon et al. 2022), and COMA (Foerster et al. 2018). A detailed introduction and implementation details for all baselines and environments are provided in Appendix  B.
Figure 4: Average episodic return curves in matrix-games environment (Right) and the payoff matrix used in this environment (Left).
Figure 5: Compatibility evaluation of QLLM in the MPE environments. The original mixing networks in RIIT, MASER, and QMIX are replaced by the TFCAF generated by QLLM, and the resulting performance is compared against the original implementations.
The Superiority of QLLM
Average Return and Convergence Speed.
Figure 3 and Figure 4 show the performance of various algorithms in certain scenarios. It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward GRF environment. Especially in the GRF environment, COMA almost fail to train the agents properly, which highlights the importance of accurate credit assignment in multi-agent environments. Furthermore, the incorporation of the IGM-Gating mechanism enables our algorithm to perform effectively in non-monotonic matrix game environments, outperforming methods like QMIX that are limited by rigid monotonicity constraints. In Listing 1, we present an example of a credit assignment function generated by LLMs for the pz-mpe-simple-spread task in the MPE environment. It can be seen that this function takes into account key factors such as the agents’ positions, their distances to landmarks, as well as the number of collisions, accurately assigning contributions to each agent. These verify the effectiveness and generalization ability of the credit assignment functions generated by our method.
Figure 6: The results of high-dimensional testing in MPE environment.
Performance in Higher Dimensional State Spaces.
We tested the performance of various algorithms in high-dimensional state spaces of certain scenarios. Taking the pz-mpe-simple-spread task as an example in the MPE environment, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially. In Figure 6, the results indicate that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in QLLM. The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult. However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
Listing 1: A credit assignment function example of pz-mpe-simple-spread environment generated by LLMs.
def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) -> torch.Tensor:
batchsize = agents_q.size(0)
n_agents = 6
global_state_reshaped = global_state.view(batchsize, n_agents, 36)
# Process landmark distances
landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2)
landmark_distances = torch.norm(landmarks_relative, dim=-1)
min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1)
pos_weights = 1.0 / min_dist_per_agent
# Process collision counts
positions = global_state_reshaped[:, :, 2:4]
diffs = positions.unsqueeze(2) - positions.unsqueeze(1)
pairwise_dist = torch.norm(diffs, dim=-1)
collision_mask = (pairwise_dist <= 0.3) & ~torch.eye(n_agents, device=positions.device).bool().unsqueeze(0)
collision_weights = collision_mask.sum(dim=-1).float()
# Combine weights and softmax
combined_weights = pos_weights - collision_weights
softmax_weights = torch.softmax(combined_weights, dim=-1)
global_q = (agents_q * softmax_weights).sum(dim=-1, keepdim=True)
return global_q
Compatibility Test.
QLLM is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values. We used the TFCAF generated by QLLM to replace the mixing networks in three value-decomposition algorithms—RIIT, MASER, and QMIX. We then conducted experiments in the MPE environment to examine how integrating TFCAF affects the performance of these value-decomposition MARL methods. The experimental results tested in MPE environments are shown in Figure 5. Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
Conclusion
In this study, we propose QLLM, a method that uses LLMs to address the credit assignment problem in multi-agent reinforcement learning through coder-evaluator framework. This approach generates training-free, task-specific credit assignment functions, reducing hallucination and improving task understanding. Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-based algorithms. Future work will explore broader applications and deployment in real-world multi-robot systems.
Appendix A LLM Prompts and Responses
Below are the prompt template, example task information, and LLM response in our work.
Role Prompt Coder’s ROLE Prompt: You are proficient in understanding tasks and writing Python code. You should fully understand the multi-agent reinforcement learning tasks and environment, including the agents, action space, observation space, and global state space provided by the user. Then, based on your understanding of the task and objectives, you need to write a function for multi-agent credit assignment, called QLLMNetwork, which is used to measure each agents’ contribution to the collective reward. The function input will be the individual Q-values of the agents and the global state, and the output will be the global Q-value. Both input and output should be torch tensors.Individual Q-values represent each agent’s expected contribution to the team’s future rewards, while the global Q-value denotes the expected overall reward of the team as a whole. The input Q-values will have a shape of torch.Size([batchsize, n_agent]), the input global state will have a shape of torch.Size([batchsize, state_dim]), and the output global Q-value should have a shape of torch.Size([batchsize, 1]).You must strictly follow the input and output dimension requirements! In the credit assignment function you wrote,You need to extract the individual components of the local observation vectors of the individual agents. You must accurately understand the meaning of each component in the global state space, and based on that, determine the weights for each Q-value. You must understand the information and suggestions provided by the user in real-time to optimize the function you design, but strictly adhere to the function’s input and output requirements! Note: 1.It must be a function, not a class. The parameters in the function are fixed and do not require training to calculate an accurate global Q-value. Do not use trainable layers such as $n$.Linear. 2.You must accurately understand the meaning of each component in the global state space, without making any assumptions or guesses! 3.When calculating the weights, do not divide by values that may be close to zero. Be careful to check intermediate results, weights for invalid numbers!This is not solved just by adding a small number to the dividend. 4.The weights are usually a combination of several components, and when calculating the individual components of the weights, it is important to pay attention to the ratio between the components and the range of the value. At the same time, the weights can be output through a softmax layer if it is necessary. 5.All tensor data should be computed on the GPU. Please write accurate code and strictly follow the template output below, without including any other explanatory text or code run sample. def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor:
# You design this place
return global_q
Evaluator’s ROLE Prompt: You are an excellent evaluator of multi-agent reinforcement learning performance. I will attempt different functions in multi-agent reinforcement learning to solve the credit assignment problem. This function is required to compute the global Q-values accurately without training. It must not use trainable layers such as $n$.Linear. Your task is to propose modifications to the function in brief plain text suggestions or return function selection results in an array format. Your suggestions must be in line with the task description and requirements, and should not be based on random assumptions or unrealistic scenarios. Please listen carefully to my subsequent instructions and do not provide additional answers! When you output plain text suggestions, please ensure that you do not output any selections. When you output selections in an array format, please ensure that you do not output plain text suggestions. Whether you output plain text suggestions or selections, consider if the function meets all the details and considerations specified in the task description. Do not use any information that has not been provided to you to answer!
Task Prompt Task Prompt of pz-mpe-simple-spread-v3: 1. Composition of the reward: We are playing Cooperative Navigation. In this task, 3 agents are asked to reach 3 landmarks. Each agent should move to get close to one landmark and avoid collision(distance<<=0.3) with other agents. The agent-team are rewarded based on how far the closest agent is to each landmark (the reward equals the negative value of the sum of the minimum distances). At the same time, the agent-team are penalized if they collide each other (-1 for each collision). At each step, each agent receives an observation array. 2. State space: This observation is represented by an array with 18 dimensions: concat([agent’s velocity, agent’s location, 3 landmarks’ relative locations, 2 other agents’ relative locations, communication code]). The agent’s velocity is 2-dimensional, including x and y components. The agent’s location is 2-dimensional, including x and y coordinates. The landmarks’ relative locations are 3x2 dimensions, including relative x and y coordinates (landmark_i.x-agent.x, landmark_i.y-agent.y). The other agents’ relative locations are 2x2 dimensions, including relative x and y coordinates (agent_i.x-agent.x, agent_i.y-agent.y). The communication code are 4 dimensions, they have no role in credit assignment. The global state space is concatenate together from the observation space of the three agents, which are 18x3 dimensions. 3. Action space: [move_left, move_right, move_down, move_up] Task Prompt of lbforaging:Foraging-8x8-2p-2f-coop-v3: The Level-Based Foraging environment focus on the coordination of involved agents. The task for each agent is to navigate the grid-world map and collect items. Each agent and food item is assigned a level and food items are randomly scattered in the environment. In order to collect a food item, agents have to choose a pick-up action next to the food item. However, such collection is only successful if 2 agents participate in the pickup at the same time. At the same time the agents involved in the pickup must be one square above, below, left and right of the food item. Agents receive team reward equal to the level of the collected food item. In each step, every agent can choose to act by moving in one of four directions or attempt to pick up a food item. The task is an 8*8 grid-world with 2 agents and 2 food items scattered in the grid. Below, we will provide additional details to global states, actions and rewards for the Level-Based Foraging environment. 1. State space: Global state vector consists of 3 component(agents component,food items component and last step action)(Total dimension: 2x3+2x3+2=14 dims) 1.Agents Component (2 agents × 3 features = 6 dims) Feature 0: agent_i’s x coordinate (i=0,1) Feature 1: agent_i’s y coordinate Feature 2: agent_i’s level 2.Food Items Component (2 Food Items × 3 features = 6 dims) Feature 0: food_i’s x coordinate (i=0,1) Feature 1: food_i’s y coordinate Feature 2: food_i’s level When a food item has been picked up, then the respective values are replaced with (-1, -1, 0).You have to be careful to handle this particular situation properly. 3.The actions just executed by the agents. (2 dims) {0,1,2…5} respectively represent {Noop, Move North, Move South, Move West, Move East, Pickup} This information is set to [-1 -1] when the environment is in the first step of each episode. 2. Action space: In Level-Based Foraging environments, each agent has six possible discrete actions to choose at each timestep: Ai = {Noop, Move North, Move South, Move West, Move East, Pickup} While the first action corresponds to the action simply staying within its grid and the last action being used to pickup nearby food, the remaining four actions encode discrete 2D navigation. Upon choosing these actions, the agent moves a single cell in the chosen direction within the grid. Picking up is only valid if the sum of the levels of the agents doing the picking up action is greater than or equal to the level of the food. At the same time, the agents involved in the pickup must be positioned one cell above, below, to the left, or to the right of the food. 3. Composition of the reward: Agents only receive non-zero team reward for picking up food items within the Level-Based Foraging environment. The reward for picking up a food item equals to the food level. Task Prompt of matrixgames:climbing-nostate-v0: This task is a two-player cooperative matrix games with a given 3×3 matrix: [[11, -30, 0], [-30, 7, 0], [0, 6, 5]] 1. Action space: The environment consists of two agents, agent_0 and agent_1, each having three possible actions(0,1,2) corresponding to the row and column indices of the matrix. 2. Composition of the reward: Once both agents select their actions, they receive a shared reward determined by the corresponding entry in the matrix. There is no communication between the two agents, and the objective is to maximize the shared reward. Be sure to prevent the strategy from falling into a local optimum, and only get the highest reward if both intelligences choose action 0. 3. State space: The state vector is 2 dims, storing the actions that the two agents have just executed. At the first step at the beginning of an episode, the state vector is [-1,-1], because the agents have not yet started executing actions. Please take care to handle this special case. Task Prompt of academy_counterattack_easy: 1. Action space: This is an offensive football game where our team controls four reinforcement learning agents, including one goalkeeper and three attacking players. The opposing team consists of one goalkeeper and one defender. The objective is to move the ball as close as possible to the opponent’s goal, located at coordinate (1.0, 0.0), through passing and off-ball movement, and to score a goal. 2. Composition of the reward: The agents receive rewards based on two components: progress reward and goal-scoring reward. The progress reward is given when a player on our team carries the ball closer to the opponent’s goal, providing a team reward. This reward is only available when our player is in possession of the ball and the ball’s distance to the opponent’s goal is greater than 0.19 and less than 0.99. The player in possession is usually the one closest to the ball. The goal-scoring reward is provided when our team successfully scores. In addition to this reward, an extra reward is given based on the distance from which the goal is scored. The farther the distance, the higher the extra reward. 3. State space: Each agent has a local observation vector of 115 dimensions: 0–7: positions of our four players, each represented by (x, y) 8–15: directions of our four players, each represented by (x, y) 16–19: positions of two opponent players, each represented by (x, y) 20–23: directions of two opponent players, each represented by (x, y) 24–87: invalid information 88–90: ball position (x, y, z) 91–93: ball direction (x, y, z) 94–96: ball possession, encoded as one-hot vectors: our team [0, 1, 0], opponent [0, 0, 1], no possession [1, 0, 0] 97–114: invalid information The global state vector is constructed by concatenating the local observations of all four agents, resulting in a total of 460 dimensions.
TFCAF template import torch import torch.nn as $n$ import torch.nn.functional as F import numpy as np class QLLMMixer($n$.Module): def __init__(self): super(QLLMMixer, self).__init__() def forward(self, agents_q, global_state): agents_q = torch.squeeze(agents_q) a = agents_q.shape[0] b = agents_q.shape[1] agents_q = agents_q.reshape(-1, agents_q.shape[-1]) global_state = global_state.reshape(-1, global_state.shape[-1]) #This is generated by LLM return (global_qagents_q.shape[-1]).reshape(a, b, 1).cuda()
LLM Generated Responses LLM Response of pz-mpe-simple-spread: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batchsize = agents_q.size(0) n_agents = 6 global_state_reshaped = global_state.view(batchsize, n_agents, 36) # Process landmark distances landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2) landmark_distances = torch.norm(landmarks_relative, dim=-1) min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1) pos_weights = 1.0 / min_dist_per_agent # Process collision counts positions = global_state_reshaped[:, :, 2:4] diffs = positions.unsqueeze(2) - positions.unsqueeze(1) pairwise_dist = torch.norm(diffs, dim=-1) collision_mask = (pairwise_dist <<= 0.3)  &  torch.eye(n_agents, device=positions.device).bool().unsqueeze(0) collision_weights = collision_mask.sum(dim=-1).float() # Combine weights and softmax combined_weights = pos_weights - collision_weights softmax_weights = torch.softmax(combined_weights, dim=-1) global_q = (agents_q  softmax_weights).sum(dim=-1, keepdim=True) return global_q LLM Response of lbforaging:Foraging-8x8-2p-2f-coop-v3: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) agents_data = global_state[:, :6].view(batch_size, 2, 3) food_data = global_state[:, 6:12].view(batch_size, 2, 3) agents_level = agents_data[:, :, 2] agents_pos = agents_data[:, :, :2].unsqueeze(2) food_pos = food_data[:, :, :2].unsqueeze(1) dxdy = torch.abs(agents_pos - food_pos).sum(dim=-1) adjacent = (dxdy == 1).float() active_food_mask = ((food_data[:, :, 0] != -1) & (food_data[:, :, 1] != -1)).unsqueeze(1) adjacent = adjacent  active_food_mask total_contribution = torch.zeros_like(agents_level) for j in range(2): adjacent_j = adjacent[:, :, j] food_level_j = food_data[:, j, 2] sum_adj_levels = (agents_level  adjacent_j).sum(dim=1) num_adj = adjacent_j.sum(dim=1) valid = ((num_adj >>= 2) & (sum_adj_levels >>= food_level_j)).float().unsqueeze(1) contribution_j = adjacent_j  food_level_j.unsqueeze(1)  valid total_contribution += contribution_j agent_contrib = agents_level  total_contribution weights = torch.softmax(agent_contrib, dim=1) global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of academy_counterattack_easy: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) n_agents = agents_q.size(1) # Extract ball information from first agent’s observation ball_pos = global_state[:, 88:90] holder_info = global_state[:, 94:97] is_my_team = holder_info[:, 1].bool() # Calculate distance to goal goal_pos = torch.tensor([1.0, 0.0], device=agents_q.device) ball_goal_dist = torch.norm(ball_pos - goal_pos, dim=1) valid_advance = is_my_team & (ball_goal_dist >> 0.19) & (ball_goal_dist << 0.99) # Calculate agent-ball distances and holder mask all_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-ball_pos[:,0], pos_y-ball_pos[:,1]], dim=1), dim=1) all_dists.append(dist.unsqueeze(1)) dist_matrix = torch.cat(all_dists, dim=1) holder_mask = F.one_hot(torch.argmin(dist_matrix, dim=1), num_classes=n_agents).bool() # Calculate agent-goal distances goal_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-goal_pos[0], pos_y-goal_pos[1]], dim=1), dim=1) goal_dists.append(dist.unsqueeze(1)) goal_dist_matrix = torch.cat(goal_dists, dim=1) # Calculate weights components advance_component = holder_mask.float()  valid_advance.unsqueeze(1)  5.0 position_component = 1.0 / (goal_dist_matrix + 1e-6) # Combine components with softmax combined = torch.where(valid_advance.unsqueeze(1), advance_component, position_component) weights = F.softmax(combined, dim=1) # Compute global Q-value global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of matrixgames:climbing-nostate-v0: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: device = agents_q.device batch_size = agents_q.shape[0] # Initialize base weights and reward matrix weights = torch.full((batch_size, 2), 0.5, device=device) reward_matrix = torch.tensor([[11, -30, 0], [-30, 7, 0], [0, 6, 5]], device=device, dtype=torch.float32) # Create mask for valid states (non-initial) valid_mask = (global_state[:, 0] >>= 0) & (global_state[:, 1] >>= 0) if valid_mask.any(): prev_actions = global_state[valid_mask].long() a0 = prev_actions[:, 0] a1 = prev_actions[:, 1] # Calculate previous rewards and optimal condition prev_rewards = reward_matrix[a0, a1] optimal_mask = (prev_rewards == 11) # Calculate deviation from optimal for each agent agent0_deviation = (a0 != 0).float() agent1_deviation = (a1 != 0).float() # Create penalty weights based on deviation penalty_weights = torch.stack([1.0 - agent0_deviation  0.5,1.0 - agent1_deviation  0.5], dim=1) # Apply optimal condition override penalty_weights[optimal_mask] = 0.5 # Maintain equal weights for optimal case # Normalize weights to maintain sum=1 weight_sums = penalty_weights.sum(dim=1, keepdim=True) normalized_weights = penalty_weights / weight_sums # Update weights for valid samples weights[valid_mask] = normalized_weights # Compute global Q-value with dynamic weighting global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q
Appendix B Details of Environments and Baselines
Baseline Algorithms
QMIX (Rashid et al. 2020b) is a representative value decomposition method designed for cooperative MARL. It factorizes the joint action-value function into individual agent utilities through a mixing network that enforces a monotonic constraint, ensuring that maximizing local utilities also maximizes the global return. The mixing network parameters are generated by a hypernetwork conditioned on the global state, enabling state-dependent flexible mixing.
QPLEX (Wang et al. 2020) builds on QMIX by introducing a duplex dueling architecture to overcome the limited expressiveness caused by strict monotonicity. It splits the value factorization into a main (advantage) stream and a complementary stream, allowing more flexible interactions while maintaining decentralized policies. This design expands the class of joint value functions that can be represented under the CTDE paradigm.
Weighted QMIX (Rashid et al. 2020a) further generalizes QMIX by learning an additional weighting factor for each agent’s contribution to the global Q-value. By relaxing the monotonicity assumption through these learnable weights, it improves the accuracy of credit assignment in tasks with more complex cooperative dependencies.
Qatten (Yang et al. 2020) enhances value decomposition by embedding a multi-head attention mechanism into the mixing network. This mechanism adaptively adjusts the importance of each agent’s local utility based on the global state context, enabling the model to focus on the most relevant agents or features when computing the global Q-value.
RIIT (Hu et al. 2021) tackles the multi-agent credit assignment problem by modeling implicit influence relationships among agents. It proposes a reward interpolation technique that redistributes the global team reward across agents, considering both immediate and long-term interdependencies, which helps reduce misattribution and improves coordination performance.
COMA (Foerster et al. 2018) is a centralized actor-critic method designed for cooperative MARL, with a focus on addressing the multi-agent credit assignment problem. It employs a centralized critic that estimates a counterfactual baseline for each agent by marginalizing out its action while keeping others fixed, enabling the computation of an advantage function that isolates each agent’s contribution. This counterfactual reasoning helps reduce variance in the policy gradient and promotes coordinated behaviors.
MASER (Jeon et al. 2022) is a value-decomposition-based MARL algorithm designed to improve credit assignment under sparse and delayed rewards. It introduces subgoals automatically generated from the experience replay buffer, which are used to construct intrinsic rewards that guide agents toward meaningful intermediate states. By augmenting standard Q-learning with subgoal-conditioned value estimation, MASER enhances exploration efficiency and coordination while remaining compatible with centralized training and decentralized execution.
Code Repository
The baseline algorithm in this experiment is implemented based on the epymarl [^5] https://github.com/uoe-agents/epymarl and pymarl2666https://github.com/hijkzzz/pymarl2 codebases. We tested MPE, LBF, and matrix games using the epymarl library, while Google Research Football was evaluated with pymarl2. Although the SMAC [^7] https://github.com/oxwhirl/smac environment is a popular benchmark for multi-agent reinforcement learning and is built upon PySC2888https://github.com/google-deepmind/pysc2, it lacks comprehensive documentation regarding the specific unit types, abilities, and map configurations used in each scenario. This obscurity hinders reproducibility and limits fine-grained control over experimental variables. Consequently, we opted for several environments with better transparency and configurability to ensure experimental rigor and facilitate controlled evaluations.
Tasks Details
Matrix Games
The cooperative matrix games environment is a simple benchmark setting in multi-agent reinforcement learning, where two or more agents simultaneously select actions from predefined action spaces. The environment is defined by a payoff matrix that specifies the global reward received based on the joint actions of all agents (below is the payoff matrix used in test environment). Cooperation is essential, as the optimal strategy often depends on coordinated decision-making among agents. In order to make sense of the global state space of this task, we include the actions executed by the agents last step.
$\begin{bmatrix}0&6&5\\ -30&7&0\\ 11&-30&0\end{bmatrix}$
Figure 7: Environment visualizations of Level-Based Foraging (LBF), Google Research Football (GRF), and Multi-Agent Particle Environments (MPE)
Level-Based Foraging
The Level-Based Foraging environment consists of tasks focusing on the coordination of involved agents. The task for each agent is to navigate the grid-world map and collect items. Each agent and item is assigned a level and items are randomly scattered in the environment. In order to collect an item, agents have to choose a certain action next to the item. However, such collection is only successful if the sum of involved agents’ levels is equal or greater than the item level. Agents receive reward equal to the level of the collected item. The tasks that were selected are denoted first by the grid-world size (e.g. 15 × 15 means a 15 by 15 grid-world). Then, the number of agents is shown (e.g. “4p” means four agents/players), and the number of items scattered in the grid (e.g. “3f” means three food items). We also have some special flags, the “2s” which denotes partial observability with a range of two squares. In those tasks, agents can only observe items or other agents as long as they are located in a square of size 5x5 centred around them. Finally, the flag “c” means a cooperative-only variant were all items can only be picked up if all agents in the level attempt to load it simultaneously. Since the global state space of the original environment is stitched together from the observation vectors of individual agents, it contains a lot of duplicated information and some environmental information is lost. Therefore, we redesigned a global state space which contains all the environment information while reducing the size of the dimension. Specifically, it includes the coordinates and levels of individual intelligences and things, as well as the actions executed by each agent last step.
Table 1: Details of LBF tasks
LBF tasks n_agents state shape action num
Foraging-2s-15x15-4p-3f 4 25 6
Foraging-2s-8x8-2p-2f-coop 2 14 6
Foraging-2s-10x10-3p-3f 3 21 6
Multi-Agent Particle Environments
Simple-Spread: In this task, three agents are trained to move to three landmarks while avoiding collisions with each other. All agents receive their velocity, position, relative position to all other agents and landmarks. The action space of each agent contains five discrete movement actions. Agents are rewarded with the sum of negative minimum distances from each landmark to any agent and a additional term is added to punish collisions among agents.
Simple-Adversary: In this task, two cooperating agents compete with a third adversary agent. There are two landmarks out of which one is randomly selected to be the goal landmark. Cooperative agents receive their relative position to the goal as well as relative position to all other agents and landmarks as observations. However, the adversary agent observes all relative positions without receiving information about the goal landmark. All agents have five discrete movement actions. Agents are rewarded with the negative minimum distance to the goal while the cooperative agents are additionally rewarded for the distance of the adversary agent to the goal landmark. Therefore, the cooperative agents have to move to both landmarks to avoid the adversary from identifying which landmark is the goal and reaching it as well. For this competitive scenario, we use a fully cooperative version where the adversary agent is controlled by a pretrained model.
Simple-Tag: In this task, three cooperating predators hunt a forth agent controlling a faster prey. Two landmarks are placed in the environment as obstacles. All agents receive their own velocity and position as well as relative positions to all other landmarks and agents as observations. Predator agents also observe the velocity of the prey. All agents choose among five movement actions. The agent controlling the prey is punished for any collisions with predators as well as for leaving the observable environment area (to prevent it from simply running away without needing to learn to evade). Predator agents are collectively rewarded for collisions with the prey. We employ a fully cooperative version of this task with a pretrained prey agent.
Table 2: Details of MPE tasks
MPE tasks n_agents state shape action num
pz-mpe-simple-spread 6 216 5
pz-mpe-simple-tag 3 48 5
pz-mpe-simple-adversary 2 16 5
Google Research Football
The Google Research Football environment is a challenging reinforcement learning platform designed to simulate football matches in a highly realistic and strategic setting. Agents control individual players and must learn to cooperate, plan, and execute complex behaviors to achieve objectives such as scoring goals or defending against opponents. The environment provides various predefined scenarios as well as full 11-versus-11 matches, supporting both single-agent and multi-agent training. In the Google Research Football environment, the action space consists of a discrete set of predefined high-level actions such as moving, passing, shooting, and sprinting, allowing agents to focus on strategic decision-making rather than low-level motor control. The Simple115 observation space is a fixed-size vector with 115 features, providing essential information about the game state, including the positions and velocities of the ball and players, the directions of movement, and other key attributes necessary for learning effective policies.
Table 3: Details of GRF tasks
GRF tasks n_agents state shape action num
academy_3_vs_1_with_keeper 4 460 19
academy_counterattack_easy 4 460 19
academy_pass_and_shoot_with_keeper 3 345 19
Appendix C Additional Experiments
Compatibility Testing of Different LLMs
We tested the performance of QLLM with three mainstream LLMs, Deepseek-R1, Deepseek-V3999https://www.deepseek.com/, and ChatGPT-4o [^10] https://openai.com/, along with other algorithms in the MPE environment. As shown in Figure 8, the experimental results indicate that mainstream LLMs outperform the baseline algorithms. Moreover, as the inference capabilities of the LLMs improve, the performance of our algorithm also improves.
Figure 8: Performance comparison of QLLM using different LLMs in the GRF environment.
Tests of the Effect of the Evaluator LLM
Most reinforcement learning algorithms enhanced by LLMs use only one kind of LLM to do their job directly. Our coder-evaluator framework features two LLMs, each with its specific role, working together to generate an accurate credit assignment function. Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in LLMs. We tested the performance of QLLM across different reinforcement learning environments without using the evaluator LLM. The experimental results (Figure 9) show that, the evaluator LLM significantly improves the algorithm’s performance in most environments.
Figure 9: Performance comparison of QLLM without evaluator LLM in the GRF environment, where QLLM-C stands for QLLM that do not use the evaluator LLM
Comparison of Algorithms Training Costs
As mentioned in the main text, QLLM replaces the mixing network with a deterministic functional expression, TFCAF, resulting in a significantly smaller number of learnable parameters. Therefore, the training cost of our algorithm is reduced. Table 4 lists the number of trainable parameters for each algorithm in the environments tested in this paper. Our algorithm reduces the number of trainable parameters by approximately 12% to 37%. This reduction directly lowers the training cost and simplifies the model structure, making the algorithm more efficient and improving performance.
Table 4: Number of learnable parameters (in K) for each environment under different algorithms
Environment / Algorithm QLLM QMIX QATTEN MADDPG RIIT QPLEX Baseline Average Reduction Rate (%)
pz-mpe-simple-spread 105.221 140.422 191.114 154.246 191.175 154.762 166.744 36.88%
pz-mpe-simple-tag 102.277 112.806 129.034 127.494 132.199 128.010 125.909 18.78%
pz-mpe-simple-adversary 101.381 107.270 118.282 122.246 124.359 122.762 118.584 14.52%
Foraging-2s-15x15-4p-3f 103.174 111.815 121.963 126.727 130.536 127.372 123.283 16.30%
Foraging-2s-8x8-2p-2f-coop 101.766 106.887 116.683 122.119 124.296 122.764 118.150 13.87%
Foraging-2s-10x10-3p-3f 102.662 109.735 120.043 124.807 127.912 125.452 121.190 15.28%
climbing-nostate-v0 100.099 103.684 110.408 118.148 118.661 118.406 113.861 12.07%
academy_3_vs_1_with_keeper 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
academy_counterattack_easy 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
academy_pass_and_shoot_with_keeper 430.099 504.884 560.760 600.084 530.869 523.046 543.129 20.83%

### v5

@@
-QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?
+listing[Listing][List of Listings]
 Abstract
-Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL).
-Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value.
-Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces.
-To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs).
-Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation.
-A custom-designed coder-evaluator framework is further employed to guide the generation and verification of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference.
-Furthermore, an IGM-Gating Mechanism enables QLLM to flexibly enforce or relax the monotonicity constraint depending on task demands, covering both IGM-compliant and non-monotonic scenarios.
-Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines.
-Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios.
-The code is available at https://github.com/zhouyangjiang71-sys/QLLM.
-Introduction
-Recent advances in cooperative multi-agent reinforcement learning (MARL)(Oroojlooy and Hajinezhad 2023) have led to widespread applications in domains such as assembly lines (Kaven et al.
-2024), logistics distribution (Yan et al.
-2022), autonomous driving (Zhang et al.
-2024b), and swarm robotics (Orr and Dutta 2023).
+Credit assignment remains a fundamental challenge in multi-agent reinforcement learning (MARL) and is commonly addressed through value decomposition under the centralized training with decentralized execution (CTDE) paradigm.
+However, existing value decomposition methods typically rely on predefined mixing networks that require additional training, often leading to imprecise credit attribution and limited interpretability.
+We propose QLLM, a novel framework that leverages large language models (LLMs) to construct training-free credit assignment functions (TFCAFs), where the TFCAFs are nonlinear with respect to the global state and offer enhanced interpretability while introducing no extra learnable parameters.
+A coder-evaluator framework is employed to ensure the correctness and executability of the generated code.
+Extensive experiments on standard MARL benchmarks demonstrate that QLLM consistently outperforms baselines while requiring fewer learnable parameters.
+Furthermore, it demonstrates generalization across a broad set of value decomposition algorithms.
+Code is available at https://github.com/MaoMaoLYJ/pymarl-qllm.
+1 Introduction
+Recent advances in cooperative multi-agent reinforcement learning (MARL) [28] have led to widespread applications in domains such as assembly lines [14], logistics distribution [42], autonomous driving [47], and swarm robotics [29].
 In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks.
-Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training.
-A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called ”lazy agents” (Liu et al.
-2023).
-To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community.
-Attention-based credit assignment methods (Soydaner 2022) incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions.
-By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions.
-This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
-Figure 1: A comparison between the traditional value decomposition method using a mixing network (Left) and our proposed novel paradigm QLLM leveraging LLMs (Right).
-The traditional approach employs a neural network to model the nonlinear relationship between local Q-values and the global Q-value, whereas QLLM capitalizes on the extensive knowledge encoded within LLMs to directly generate a training-free credit assignment function.
-Policy gradient approaches (Agarwal et al.
-2021) tackle credit assignment from a differentiable optimization perspective.
-Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability.
-These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation.
-In addition, value decomposition methods (Sunehag et al.
-2017) have become increasingly popular.
-These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent.
-The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution.
-These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
-However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for each agents.
-A representative example of this is the mixing network employed in value decomposition methods.
-These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability (Linardatos, Papastefanopoulos, and Kotsiantis 2020).
-With the advancement of large language models (LLMs) (Kasneci et al.
-2023), new paradigms have emerged to address many long-standing challenges in MARL.
-By leveraging extensive pre-trained knowledge (Petroni et al.
-2019), superior generalization capacity (Anil et al.
-2022), and powerful code generation abilities (Xu et al.
-2022a), LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines.
-Recent studies (Qu et al.
-2025) have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals.
-Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines.
-Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency (Sanh et al.
-2021).
-In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality (Zhang et al.
-2023b).
-Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
-In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network.
-To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation.
-Within this framework, LLMs act as both a code generator and a code evaluator (Desmond et al.
-2024).
-Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT (Rashid et al.
-2020b; Su, Adams, and Beling 2021; Hu et al.
-2021).
-Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability.
-Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues (Martino, Iannelli, and Truong 2023) commonly encountered by LLMs.
-This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network.
-Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks.
-An IGM-Gating Mechanism is further introduced to flexibly switch the monotonicity constraint on or off, adapting to tasks where the Individual-Global-Max principle holds or does not.
+Since agents are typically trained using a shared team-level reward, accurately attributing individual contributions is critical for effective policy optimization [27].
+This challenge is commonly referred to as the credit assignment problem, where inaccurate attribution can lead to suboptimal coordination behaviors, such as the emergence of lazy agents [22].
+To address the credit assignment problem in multi-agent systems, a variety of approaches have been explored in the literature.
+Policy gradient approaches [1, 10, 6] address credit assignment through a centralized critic, where joint value or advantage functions are used to evaluate agents’ actions under shared rewards, and individual policies are updated based on agent-specific gradients derived from these centralized evaluations.
+In addtion, value decomposition methods [38] have emerged as a dominant paradigm under the centralized training with decentralized execution (CTDE) framework [23].
+These methods construct a joint value function that is decomposed into individual utility functions, ensuring that the optimal joint policy can be recovered from local components while enabling scalable and coordinated training.
+In practice, value decomposition methods typically rely on neural-network-based mixing networks [35] to parameterize the credit assignment process.
+However, similar to other neural credit assignment approaches, they require additional training, incur substantial optimization overhead, and often lack semantic interpretability [21].
+With the rapid advancement of large language models (LLMs) [13], they have demonstrated strong capabilities in natural language processing (NLP), including text understanding [2] and code generation [12].
+Building on these capabilities, LLMs have increasingly been explored as general-purpose components in reinforcement learning (RL).
+In this context, LLMs can support RL in multiple ways, such as assisting reward design [25], constructing abstract world models [31], or directly serving as decision-making policies [18].
+By injecting structured prior knowledge and high-level reasoning into the learning process, LLMs complement conventional data-driven optimization in complex environments.
+Figure 1: Comparison between the traditional value decomposition method using the mixing network (Left) and our proposed novel paradigm QLLM leveraging LLMs (Right).
+Motivated by these developments, we investigate the use of LLMs to address the credit assignment problem in multi-agent systems.
+This paper introduces QLLM, a novel value decomposition framework that eliminates the need for a learned mixing network, as shown in Figure 1.
+To this end, a general coder-evaluator framework is proposed to support reliable LLM-based code generation.
+Within this framework, LLMs assume dual roles as both a code generator and a code evaluator [5], enabling the synthesis and validation of executable functions.
+Given task specifications and code generation instructions, the code generator produces a nonlinear Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in value decomposition algorithms.
+Because TFCAF is generated directly by LLMs, it requires no additional training and offers improved semantic interpretability, while remaining robust to complex state representations.
+To further enhance reliability, a code verification mechanism is incorporated to ensure the correctness and executability of the generated code, thereby mitigating hallucination issues commonly observed in LLM-based code generation [26].
+As a result, the proposed framework enables the construction of a dependable TFCAF without direct interaction with the environment, which can serve as a drop-in replacement for the mixing network and be readily integrated into a broad class of value decomposition methods.
 The main contributions of this study are summarized as follows:
   1. 1.
-    We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF.
-This framework is generalizable and can be extended to other MARL scenarios enhanced by LLMs.
+    A coder-evaluator framework is adopted to enable the zero-shot construction of a reliable, training-free credit assignment function, termed TFCAF, via LLM-based code generation to the credit assignment problem in MARL.
   2. 2.
-    QLLM is proposed by combining TFCAF and value function decomposition while remaining compatible with most MARL algorithms that employ mixing networks for credit assignment.
-Unlike existing methods, QLLM can accurately infer each agent’s contribution without interacting with the environment.
+    A novel value decomposition framework, QLLM, is introduced by integrating TFCAF into MARL algorithms based on mixing network, allowing credit assignment without additional training or environment interaction.
   3. 3.
-    Extensive experiments demonstrate that QLLM outperforms existing credit assignment algorithms, particularly in scenarios involving high-dimensional state spaces.
-Moreover, the integration of LLMs endows the contribution assignment among agents with semantic interpretability.
-Related Work
-Credit Assignment in MARL
+    Extensive experiments demonstrate that QLLM consistently outperforms existing credit assignment methods and provides improved semantic interpretability, particularly in complex state settings.
+2 Related Work
+2.1 Credit Assignment in MARL
 In recent years, the credit assignment problem has emerged as a fundamental challenge in multi-agent systems.
-And policy-based approaches (Yu et al.
-2022; Zhang et al.
-2024a) were among the first to introduce credit assignment mechanisms.
-MADDPG (Lowe et al.
-2017) extends the DDPG (Lillicrap et al.
-2015) framework to multi-agent settings by incorporating centralized training with decentralized execution.
-This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments.
-MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios.
-Similarly, COMA (Foerster et al.
-2018) leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG.
-MAAC (Iqbal and Sha 2019) further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment.
+Policy-based approaches [44, 45] were among the first to introduce credit assignment mechanisms.
+MADDPG [23] extends the DDPG [19] framework to multi-agent settings by incorporating CTDE.
+This approach utilizes an actor-critic architecture [15] with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios.
+Similarly, COMA [6] adopts a centralized critic and introduces a counterfactual baseline to estimate agent-specific advantage functions.
+MAAC [10] further enhances policy-based methods by introducing an attention mechanism into the critic, leading to more precise credit assignment.
 However, most of these algorithms are not directly applicable to tasks with shared rewards.
-Another widely adopted paradigm is the value decomposition approach (Rashid et al.
-2020b; Xu et al.
-2022b), which explicitly decomposes the globally shared return.
-As one of the earliest methods in this category, VDN (Sunehag et al.
-2017) models the global value function as a linear summation of individual value functions.
+Another widely adopted paradigm is the value decomposition approach [35, 41], which explicitly decomposes the globally shared return.
+As one of the earliest methods in this category, VDN [38] models the global value function as a linear summation of individual value functions.
 While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition.
-QMIX (Rashid et al.
-2020b) improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations.
+QMIX [35] improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations.
 However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments.
-QTRAN (Son et al.
-2019) and QPLEX (Wang et al.
-2020) mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity.
-LICA (Zhou et al.
-2020) adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information.
-Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks.
-RIIT (Hu et al.
-2021) goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies.
-Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value.
-Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training.
-Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
-LLM-enhanced Reinforcement Learning
-Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning (Zhang et al.
-2023a; Li et al.
-2024).
-In the SMAC-R1 framework (Deng et al.
-2024), agents utilize LLMs to generate decision tree code by providing task descriptions.
-The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards.
-Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments.
-They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes.
-For instance, the Reflexion framework (Shinn et al.
-2023) introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters.
-Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions.
-Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks.
-The EUREKA framework (Ma et al.
-2023) achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
-These examples demonstrate the powerful information processing and code generation capabilities of LLMs.
-Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL.
+QTRAN [37] and QPLEX [40] mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity.
+RIIT [9] goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies.
+Nevertheless, RIIT ultimately retains a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value.
+While the mixing network plays a central role in implementing credit assignment, it often suffers from limited interpretability and slow convergence during training.
+As a result, despite the success of value decomposition methods, achieving efficient, precise, and interpretable credit assignment in multi-agent systems remains an open challenge.
+2.2 LLM-Enhanced RL
+Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into RL to enhance performance in multitask learning, sample efficiency, and high-level task planning [46, 17].
+Recent studies have explored the use of LLMs to generate structured decision policies or programmatic representations from task descriptions, enabling more interpretable and adaptable control strategies [4, 18].
+In addition, LLMs can provide natural language rationales for agent behaviors, improving the transparency and explainability of decision-making processes [3, 32].
+LLMs have been investigated as tools for shaping reward signals, where automatically generated reward functions or feedback mechanisms help guide agents toward more effective behaviors [25, 24].
+Collectively, these works highlight the potential of LLMs as flexible components for enhancing learning signals and structural priors in RL.
+Recent work has begun to explore the use of LLMs for temporal and agent-level credit assignment in MARL.
 Qu et al.
-(2025) were the first to introduce LLMs into the credit assignment problem.
-They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly.
-This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms.
-However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components.
-Additionally, Lin et al.
-(2025) employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions.
-These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance.
-Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
-Preliminaries
-Decentralized Partially Observable Markov Decision Process
-In this work, we focus on a fully cooperative MARL setting that can be formalized as a decentralized partially observable Markov decision process (Dec-POMDP).
+[33] proposed latent rewards to capture cooperative relationships, effectively addressing credit assignment under sparse rewards and outperforming existing baselines.
+Meanwhile, Lin et al.
+[20] employed LLMs to learn potential-based reward functions across multiple queries, allowing for more accurate estimation of individual agent contributions.
+Despite their effectiveness, these methods primarily focus on reward decomposition and remain limited to settings compatible with single-agent reinforcement learning formulations, which restricts their scalability and applicability to general MARL algorithms.
+To overcome these limitations, QLLM is introduced as a multi-agent credit assignment framework that uses LLM-based code generation to construct training-free credit assignment functions compatible with value decomposition methods.
+3 Preliminaries
+3.1 Dec-POMDP
+In this work, we focus on a fully cooperative MARL setting formalized as a decentralized partially observable Markov decision process (Dec-POMDP[27].
 A Dec-POMDP is typically defined by a tuple $G=\langle S,A,P,r,Z,O,n,\gamma\rangle$.
-Here, $n$ agents indexed by $i\in\{1,\dots,n\}$ operate in an environment with true underlying state $s\in S$.
+In this setting, $n$ agents indexed by $i\in\{1,\dots,n\}$ operate in an environment with a true underlying state $s\in S$.
 At each timestep, agent $i$ selects an action $a^{i}\in A$ based solely on its private observation $z_{i}\in Z$, which is sampled according to the observation function $O(s,i)$.
 The joint action of all agents is denoted by $\boldsymbol{a}\in\boldsymbol{A}$.
-The environment evolves according to the transition probability function $P(s^{\prime}\mid s,\boldsymbol{a}):S\times\boldsymbol{A}\to[0,1]$, which specifies the likelihood of transitioning to state $s^{\prime}$ given the current state $s$ and joint action $\boldsymbol{a}$.
-All agents cooperate to maximize a shared reward signal defined by the function $r(s,\boldsymbol{a}):S\times\boldsymbol{A}\to\mathbb{R}$.
-The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward $\sum_{j=0}^{\infty}\gamma^{j}r_{t+j}$, where $\gamma\in[0,1)$ is a discount factor determining the weight of future rewards.
-A key challenge in Dec-POMDPs lies in the credit assignment problem.
-This issue arises because all agents within the system share a common reward structure, making it inherently difficult to accurately attribute individual contributions to the collective outcome.
-Furthermore, the environment is perceived as non-stationary from the viewpoint of each agent.
-This non-stationarity poses significant challenges in achieving effective coordination and optimal decision-making, as agents must adapt to the changing strategies of others while managing incomplete state representations.
-Figure 2: QLLM is designed within a coder-evaluator framework to autonomously generate high-quality, training-free credit assignment functions (TFCAFs).
-In each iteration, the coder LLM $M_{\text{coder}}$ produces $K$ candidate functions based on a task prompt and a role-specific instruction, referred to as the coder prompt.
-These candidate functions are initially assessed using environmental state information and local Q-values.
-If any runtime or syntax errors are identified, the faulty candidates are discarded and regenerated.
-Subsequently, the evaluator LLM $M_{\text{evaluator}}$ reviews the candidate functions, selects the most promising one based on its evaluator prompt This process is generated by LLMs through structured prompts without any need for human intervention.
-Value Decomposition and IGM Principle
-Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions $Q_{i}$ of individual agents into a global Q-value function $Q_{\text{tot}}$.
-The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint.
-The IGM principle (Hong, Jin, and Tang 2022) requires that the joint action that maximizes the global action-value function $Q_{\text{tot}}$ must correspond to the combination of each agent’s individually optimal action.
+The environment evolves according to the transition probability function $P(s^{\prime}\mid s,\boldsymbol{a}):S\times\boldsymbol{A}\to[0,1]$, which specifies the probability of transitioning to state $s^{\prime}$ given the current state $s$ and joint action $\boldsymbol{a}$.
+All agents cooperate to maximize a shared reward signal defined by the reward function $r(s,\boldsymbol{a}):S\times\boldsymbol{A}\to\mathbb{R}$.
+The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward $\sum_{j=0}^{\infty}\gamma^{j}r_{t+j}$, where $\gamma\in[0,1)$ is a discount factor.
+Dec-POMDPs pose inherent challenges due to partial observability and non-stationarity.
+Beyond these difficulties, a central challenge in Dec-POMDPs is the credit assignment problem, which is the primary focus of this work.
+Because agents typically optimize a shared team reward, accurately attributing individual contributions to the collective outcome is non-trivial and plays a critical role in effective coordination and learning.
+3.2 Value Decomposition and IGM Principle
+Traditional value decomposition methods employ a mixing network to aggregate the local action-value functions $Q_{i}$ of individual agents into a global action-value function $Q_{\text{tot}}$ following the Individual-Global-Maximum (IGM) principle [35],whereby the joint action that maximizes $Q_{\text{tot}}$ must correspond to the combination of each agent’s individually optimal action.
 Formally, the IGM principle can be expressed as:
 $\arg\max_{\boldsymbol{a}}Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{a})=\left(\begin{array}[]{c}\arg\max_{a^{1}}Q_{1}\left(\tau^{1},a^{1}\right)\\ \vdots\\ \arg\max_{a^{n}}Q_{n}\left(\tau^{n},a^{n}\right)\end{array}\right),$ (1)
-where $\boldsymbol{\tau}\in T^{n}$ represents the joint action-observation histories of all agents.
+where $\boldsymbol{\tau}=(\tau^{1},\dots,\tau^{n})$ represents the joint action-observation histories of all agents.
 This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality.
-Building upon this, Qatten (Yang et al.
-2020) provides a theoretical extension, showing that $Q_{tot}$ can be expressed as a linear combination of the individual $Q_{i}$ functions:
+Building upon this principle, Qatten [43] provides a theoretical extension, showing that $Q_{\text{tot}}$ can be expressed as a linear combination of the individual $Q_{i}$ functions:
 $Q_{\text{tot}}(s,\boldsymbol{a})=\sum_{i=1}^{n}\sum_{h=1}^{H}w_{i,h}(s)Q_{i}(\tau^{i},a^{i})+b.$ (2)
-Here, $w_{i,h}(s)$ denotes a coefficient that depends on the state $s$, $H$ is the number of attention heads.
+In this formulation, $w_{i,h}(s)$ denotes a coefficient that depends on the state $s$, and $H$ is the number of attention heads.
 Theoretically, these coefficients can be derived from a higher-order functional expansion of $Q_{\text{tot}}$ with respect to $Q_{i}$, where the contribution of each term diminishes rapidly with increasing derivative order.
 This indicates that the influence of higher-order interaction terms on $Q_{\text{tot}}$ is sufficiently small to be ignored or effectively absorbed into a bias term $b$, allowing for a simplified model formulation and reduced computational complexity.
-Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
-Method
-This section presents a detailed discussion of the implementation of each component of QLLM.
-We first propose the concept of the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and more relaxed joint value modeling.
-This provides a foundation for adaptively handling different types of multi-agent tasks.
-Then we introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents.
+4 Methodology
+This section presents a detailed discussion of the implementation of each component of QLLM, whose overall framework is illustrated in Figure 2.
+We first introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents.
 Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs.
-However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination (Tonmoy et al.
-2024) and a lack of rigorous reasoning.
+However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination [39] and a lack of rigorous reasoning.
 These limitations motivate the design of our proposed coder-evaluator framework.
-By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
-IGM-Gating Mechanism
-To flexibly handle diverse coordination demands across tasks, we introduce the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and relaxed joint value modeling.
-This mechanism allows for conditional enforcement of the monotonicity constraint in the credit assignment.
-Definition 1 (IGM-Gating Mechanism).
-Let $Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})$ be the joint action-value function, and $\{Q_{i}(s_{i},a_{i})\}_{i=1}^{n}$ the individual utility functions.
-Define a binary gating variable $\gamma\in\{0,1\}$, where $\gamma=1$ indicates that IGM is enforced.
-The IGM-Gating Mechanism imposes the following conditional monotonicity constraint:
-$\forall i,\quad\gamma=1\Longrightarrow\frac{\partial Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})}{\partial Q_{i}(s_{i},a_{i})}\geq 0.$ (3)
-When the gating variable is inactive ($\gamma=0$), no monotonicity constraint is enforced, allowing $Q_{\text{tot}}$ to freely capture non-monotonic dependencies.
-Training-Free Credit Assignment Function
-As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions.
-The network takes as input the individual Q-values of all agents, denoted as $\boldsymbol{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$, and produces the joint Q-value $Q_{\text{tot}}$ as output.
-To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism (Ha, Dai, and Le 2017).
-Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state $s$.
-This design facilitates state-aware adaptation while ensuring compatibility with CTDE framework, making deployment straightforward.
+By integrating these components, QLLM is able to generate highly accurate TFCAF functions for multi-agent credit assignment, providing an effective alternative to the traditional mixing networks used in value decomposition approaches.
+Figure 2: QLLM framework: $M_{\text{coder}}$ generates candidate TFCAFs, and $M_{\text{evaluator}}$ selects the optimal function after automated verification.
+4.1 Training-Free Credit Assignment Function
+Traditional mixing networks often rely on hypernetwork [8] to generate state-dependent parameters.
 However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
-To address these issues, we propose the Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM.
-This module is generated by LLMs through iterative error correction, based on a series of easily accessible task-related prompts.
+To address these issues, we propose a training-free credit assignment function (TFCAF), which serves as the core module of QLLM.
+This module is generated by LLMs based on a series of task-related prompts, with correctness ensured through structured code verification.
 As illustrated in Equation (2), the global Q-value function $Q_{\text{tot}}$ can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term.
-Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF, which we formally define below.
-Definition 2.
-A Training-Free Credit Assignment Function (TFCAF) is defined as:
-$Q_{\text{tot}}\left(s,\boldsymbol{a}\right)=\sum_{i=1}^{n}f_{w}^{i}(s)\,Q_{i}\left(\tau^{i},a^{i}\right)+f_{b}(s),$ (4)
-where $f_{w}^{i}(s)$ and $f_{b}(s)$ are general nonlinear functions of the global state $s$ generated by LLMs before training.
-As illustrated in Definition 1, when the gating signal $\gamma=1$, the function $f_{w}^{i}(s)$ is strictly constrained to be non-negative.
-In contrast, when $\gamma=0$, this constraint is lifted.
-Benefiting from the rich prior knowledge of LLMs, the appropriate value of $\gamma$ can be easily inferred from the task description, enabling the model to adapt to different types of tasks.
-For instance, in matrix game scenarios, where individual optimal decisions do not necessarily align with the global optimum, the LLM may assign a negative value to $f_{w}^{i}(s)$.
-The coefficients $f_{w}^{i}\left(s\right)$ and the bias term $f_{b}\left(s\right)$ make the overall TFCAF a nonlinear mapping, although its final combination with the local $Q_{i}$ remains linear in form.
-Both components are automatically produced by LLMs through the coder-evaluator framework, which we detail next along with the iterative correction and its role in credit assignment for multi-agent systems.
-Coder Evaluator Framework
-Due to the susceptibility of LLMs to hallucinations, the process of generating TFCAF must be designed to be robust.
-To address this challenge, we propose the coder-evaluator framework, which automatically generates TFCAF through a modular and self-correcting process, ensuring both accurate credit assignment and logical consistency in code generation.
-This framework employs two LLMs with distinct yet complementary roles: the coder LLM $M_{coder}$, responsible for generating candidate TFCAFs, and the evaluator LLM $M_{evaluator}$, which critically evaluates these candidates.
-All of these are generated by LLMs through structured prompts without any need for human intervention.
-The interaction between the two models is driven by two types of prompts: task prompts ($P_{task}$) specify the meanings of each dimension in the global state space, the action space of the agents, and the composition of the reward function.
-This information is objective and can be easily obtained from the environment; role prompts ($P_{role}$) define the fixed responsibilities and operational roles of each LLM within the framework.
-The role prompts are further categorized into coder prompts and evaluator prompts, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively.
-This clear separation and stability of prompt design make QLLM easy to adapt to new cooperative tasks without extensive manual prompt tuning or subjective rewriting (see in Appendix A).
+Motivated by this decomposition paradigm, a similar formulation is adopted in the design of TFCAF.
+Its mathematical expression is as follows:
+$Q_{\text{tot}}\left(s,\boldsymbol{a}\right)=\sum_{i=1}^{n}{f_{w}^{i}}\left(s\right)Q_{i}\left(\tau^{i},a^{i}\right)+f_{b}\left(s\right).$ (3)
+The coefficient $f_{w}^{i}\left(s\right)$ is a state-dependent weight, and $f_{b}\left(s\right)$ is a bias term that also depends on the global state.
+Both are key components of the TFCAF, which are generated by LLMs.
+A detailed theoretical proof justifying this decomposition formulation is provided in Appendix 0.D.
+In the following, we provide a detailed explanation of how the TFCAF is generated, as well as how it facilitates credit assignment in multi-agent systems.
+4.2 Coder-Evaluator Framework
+Our framework employs two distinct LLM roles: the coder LLM $M_{coder}$ for generating candidate TFCAFs and the evaluator LLM $M_{evaluator}$ for critical selection and validation, ensuring the robustness of the generated functions with minimal human involvement.
+The interaction between the two models is driven by two types of prompts.
+Specifically, Task prompts ($P_{task}$) provides environment-specific details, while Role prompts ($P_{role}$) define the fixed responsibilities and operational roles of each LLM within the framework.
+These role prompts are further categorized into coder prompt and evaluator prompt, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively.
+This separation between task-specific and role-specific prompts enables QLLM to adapt to new cooperative tasks with minimal prompt modification, without requiring extensive manual tuning or subjective rewriting.
+Detailed prompt templates used in the framework are provided in Appendix 0.A.
 The generation of TFCAF can be formally described through the following three clearly-defined steps.
 Initial Generation.
-Firstly, $M_{coder}$ generates $K$ candidate credit assignment functions by processing both the role-specific and task-specific prompts:
-$\phi_{1},\phi_{2},...,\phi_{K}\ \Leftarrow\ M_{coder}\big(P_{role},\ P_{task}\big),$ (5)
-where $\phi_{1},\phi_{2},\dots,\phi_{K}$ denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases.
+At the initial stage, the Coder LLM $M_{\text{coder}}$ generates a set of candidate credit assignment functions conditioned on the provided prompts.
+Specifically, given the role-specific prompts $P_{\text{role}}$ and the task-specific prompts $P_{\text{task}}$, $M_{\text{coder}}$ produces $K$ candidate functions:
+$\phi_{1},\phi_{2},\dots,\phi_{K}\ \Leftarrow\ M_{\text{coder}}\big(P_{\text{role}},P_{\text{task}}\big),$ (4)
+where each $\phi_{k}$ represents a candidate TFCAF.
+Each generated function explicitly specifies how to compute the state-dependent weight functions $\{f_{w}^{i}(s)\}_{i=1}^{n}$ and the bias term $f_{b}(s)$ in Equation (3), typically in the form of executable code or programmatic expressions.
+Consequently, each $\phi_{k}$ defines a complete mapping from the global state $s$ to the mixing coefficients used for credit assignment, without introducing additional learnable parameters.
 Error Detection and Correction.
-Due to the inherent hallucination tendencies of LLMs, candidate functions might contain syntax or dimensional errors.
-To address this, each candidate TFCAF is compiled and run once with any available input state and local Q-values.
-If an error occurs during execution, the error message generated by the compiler will be fed back to $M_{coder}$, prompting regeneration of the faulty TFCAF:
-$\phi_{correct}\,\,\Leftarrow\,\,M_{coder}\big(P_{role},\,\,P_{task},\ error\big).$ (6)
-This correction step repeats until all $K$ candidate TFCAFs can execute successfully without errors.
-Selection of TFCAF.
-Subsequently, $M_{evaluator}$ evaluates these validated candidate TFCAFs, selecting the most promising one based on task relevance, interpretability, and code clarity.
-Notably, the selection of the best candidate is fully guided by structured prompts and performed automatically by $M_{evaluator}$, without any manual or subjective intervention:
-$choice\,\Leftarrow\,\,M_{evaluator}\big(P_{role},\ P_{task},\ {\Phi}\big),$ (7)
-where ${\Phi}=\{\phi_{1},\phi_{2},...,\phi_{K}\}$.
-Algorithm 1 Pseudo-code for QLLM
-1:Part 1: TFCAF Generation (LLM-based)
-2: $\phi_{1},\phi_{2},\dots,\phi_{K}\leftarrow M_{\text{coder}}(P_{\text{role}},\,P_{\text{task}})$ $\triangleright$ Generate $K$ candidate functions
-3:for all $\phi_{k}\in\{\phi_{1},\dots,\phi_{K}\}$ do
-4:  if $\phi_{k}$ fails on environment samples then
-5:   $\phi_{k}\leftarrow M_{\text{coder}}(P_{\text{role}},P_{\text{task}},\text{error})$ $\triangleright$ Regenerate function based on error message
-6:  end if
-7:end for
-[…]
-[…]
-10:Generate $f_{\text{TFCAF}}$ from $\phi_{\text{final}}$ using predefined template
-11:Part 2: Agents Optimization (RL-based)
-12:for episode = 1 to $M$ do
-13:  for each agent $i$ do
-14:   $a^{i}\leftarrow\arg\max_{a}Q_{i}(\tau^{i},a;\theta)$
-15:  end for
-16:  Compute global Q-value:
-17:   $Q_{\text{tot}}(s,\boldsymbol{a})=f_{\text{TFCAF}}(Q_{1},...,Q_{n};s)$
-18:  Store transition $(s,\boldsymbol{a},r,s^{\prime})$ in buffer $\mathcal{D}$
-19:  Sample batch and compute TD target:
-20:   $y=r+\gamma\max_{\boldsymbol{a}^{\prime}}Q_{\text{tot}}(s^{\prime},\boldsymbol{a}^{\prime};\theta^{-})$
-21:  Update parameters $\theta$ by minimizing the TD loss:
-22:   $\mathcal{L}(\theta)=\mathbb{E}\left[\left(y-Q_{\text{tot}}(s,\boldsymbol{a};\theta)\right)^{2}\right].$
-23:end for
-Training Procedure
+Due to the inherent tendency of LLMs to produce hallucinated or imperfect code, the generated candidate functions may contain syntax errors, dimensional mismatches, or invalid operations.
+To ensure executability, each candidate TFCAF is first compiled and executed once using a set of available input states and corresponding local action-value estimates.
+If an execution error occurs, the corresponding error message returned by the compiler or runtime environment is collected and provided as structured feedback to the Coder LLM $M_{\text{coder}}$.
+Conditioned on this feedback, $M_{\text{coder}}$ regenerates a corrected version of the faulty TFCAF:
+$\phi_{\text{correct}}\ \Leftarrow\ M_{\text{coder}}\big(P_{\text{role}},\,P_{\text{task}},\,\text{error}\big).$ (5)
+This procedure continues until all $K$ candidate TFCAFs execute successfully, ensuring basic syntactic and execution-level correctness before they are passed to subsequent evaluation stages.
+Evaluation of TFCAF.
+After execution-level validation, the Evaluator LLM $M_{\text{evaluator}}$ evaluates the candidate TFCAFs to identify the most suitable one for credit assignment.
+Given the validated candidate set $\Phi=\{\phi_{1},\phi_{2},\dots,\phi_{K}\}$, $M_{\text{evaluator}}$ focuses on whether a TFCAF provides a coherent and task-aligned mapping from the global state to agent-wise credit weights, rather than relying on empirical performance metrics.
+The evaluation and final choice are fully guided by structured role and task prompts and are performed automatically by $M_{\text{evaluator}}$, without manual ranking or subjective human intervention:
+$\phi_{\text{choice}}\ \Leftarrow\ M_{\text{evaluator}}\big(P_{\text{role}},\,P_{\text{task}},\,\Phi\big),$ (6)
+where the selected TFCAF is subsequently adopted as the credit assignment function in QLLM.
+Subsequently, $\phi_{\text{choice}}$ is fed back into $M_{\text{coder}}$ to serve as a reference for the next cycle of generation, and QLLM iteratively repeats this process for a total of $T$ rounds to ensure continuous logic refinement.
+4.3 Training Procedure
 With the coder-evaluator framework, we obtain a training-free credit assignment function.
@@
 where $f_{\text{TFCAF}}$ is a parameterized function in which the weights are allowed to take arbitrary values, while the biases are also unconstrained.
-$f_{\text{TFCAF}}$ receives the agent-specific Q-values $Q_{i}$ as input and utilizes the global state $s$ to generate the corresponding weights and biases through the weight module $f_{w}^{i}$ and the bias module $f_{b}$, respectively.
+$f_{\text{TFCAF}}$ receives the individual action-value functions $Q_{i}$ as input and utilizes the global state $s$ to generate the corresponding weights $f_{w}^{i}$ and bias $f_{b}$ through the coder-evaluator framework, respectively, both of which are modeled as nonlinear functions of the global state.
 Notably, all parameters in the TFCAF are generated by LLMs and remain fixed during training.
@@
 $\mathcal{L}(\theta)=\mathbb{E}_{(s,\boldsymbol{a},r,s^{\prime})}\bigg[\big(y-Q_{\text{tot}}(s,\boldsymbol{a};\theta)\big)^{2}\bigg],$ (8)
-where the target value $y$ is computed using Bellman equation:
+where the target value $y$ follows the Bellman equation:
 $y=r+\gamma\max_{\boldsymbol{a}^{\prime}}Q_{\text{tot}}(s^{\prime},\boldsymbol{a}^{\prime};\theta^{-}),$ (9)
-with $\gamma$ being the discount factor. $\theta^{-}$ denotes the parameters of the target agent network.
+where $\theta^{-}$ denotes the parameters of the target agent network.
 Compared to other neural network-based approaches for multi-agent credit assignment, QLLM involves fewer trainable parameters.
@@
 Meanwhile, the incorporation of TFCAF allows QLLM to effectively exploit prior knowledge while employing an interpretable credit assignment mechanism.
-Experiments
-Figure 3: Average episodic return curves for selected tasks in GRF and MPE environments (the results for LBF are shown in Appendix C).
-To comprehensively evaluate the proposed method, the experiments were designed with six main objectives: (1) to assess the performance of the QLLM algorithm in terms of average return across multi-agent environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other mixing-based algorithms; (4) to demonstrate the compatibility of QLLM with various LLMs; (5) to validate the effectiveness of the proposed coder-evaluator framework under various parameters; (6) to demonstrate the superiority of our method in terms of training costs.
-Experimental Setups
+5 Experiments
+To comprehensively evaluate the efficacy and versatility of the proposed QLLM framework, we design and conduct experiments focused on the following objectives: (1) assessing the performance of QLLM across a diverse range of cooperative multi-agent environments; (2) validating the generality and compatibility of QLLM by integrating it with various established value decomposition algorithms; (3) evaluating its robustness to high-dimensional global state spaces by systematically increasing the number of agents and state dimensionality; and (4) investigating the interpretability of TFCAF to provide insights into the generated credit assignment logic.
+Furthermore, additional ablation studies and detailed comparisons of the number of learnable parameters across different algorithms are provided in Appendix 0.B.
+5.1 Experimental Setups
 We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms.
-The primary LLM used in our method is the DeepSeek-R1 (Guo et al.
-2025) inference model.
-For each experiment, the reported return values represent the mean over five independent training runs, with 90% confidence intervals indicated by error bars.
-Unless otherwise specified, all environment parameters are set to their default values.
-Our experimental environments include: (1) Cooperative Matrix Games [^1] https://github.com/uoe-agents/matrix-games, used to validate the algorithm’s performance in non-monotonic scenarios, which may violate the IGM principle; (2) Multi-agent Particle Environments (MPE) [^2] https://pettingzoo.farama.org/environments/mpe/, for evaluating performance in high-dimensional state spaces; (3) Level-Based Foraging (LBF) [^3] https://github.com/uoe-agents/lb-foraging, for testing in sparse reward settings; and (4) Google Research Football (GRF) [^4] https://github.com/google-research/football, to assess performance in complex, realistic environments.
-We compare QLLM with six commonly used baseline algorithms: QMIX (Rashid et al.
-2020b), QPLEX (Wang et al.
-2020), Weighted QMIX (Rashid et al.
-2020a), Qatten (Yang et al.
-2020), RIIT (Hu et al.
-2021), MASER (Jeon et al.
-2022), and COMA (Foerster et al.
-2018).
-A detailed introduction and implementation details for all baselines and environments are provided in Appendix  B.
-Figure 4: Average episodic return curves in matrix-games environment (Right) and the payoff matrix used in this environment (Left).
-Figure 5: Compatibility evaluation of QLLM in the MPE environments.
-The original mixing networks in RIIT, MASER, and QMIX are replaced by the TFCAF generated by QLLM, and the resulting performance is compared against the original implementations.
-The Superiority of QLLM
-Average Return and Convergence Speed.
-Figure 3 and Figure 4 show the performance of various algorithms in certain scenarios.
-It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward GRF environment.
-Especially in the GRF environment, COMA almost fail to train the agents properly, which highlights the importance of accurate credit assignment in multi-agent environments.
-Furthermore, the incorporation of the IGM-Gating mechanism enables our algorithm to perform effectively in non-monotonic matrix game environments, outperforming methods like QMIX that are limited by rigid monotonicity constraints.
-In Listing 1, we present an example of a credit assignment function generated by LLMs for the pz-mpe-simple-spread task in the MPE environment.
-It can be seen that this function takes into account key factors such as the agents’ positions, their distances to landmarks, as well as the number of collisions, accurately assigning contributions to each agent.
-These verify the effectiveness and generalization ability of the credit assignment functions generated by our method.
-Figure 6: The results of high-dimensional testing in MPE environment.
-Performance in Higher Dimensional State Spaces.
-We tested the performance of various algorithms in high-dimensional state spaces of certain scenarios.
-Taking the pz-mpe-simple-spread task as an example in the MPE environment, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially.
-In Figure 6, the results indicate that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in QLLM.
-The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult.
-However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
-Listing 1: A credit assignment function example of pz-mpe-simple-spread environment generated by LLMs.
-def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) -> torch.Tensor:
-batchsize = agents_q.size(0)
-n_agents = 6
-global_state_reshaped = global_state.view(batchsize, n_agents, 36)
-# Process landmark distances
-landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2)
-landmark_distances = torch.norm(landmarks_relative, dim=-1)
-min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1)
-pos_weights = 1.0 / min_dist_per_agent
-# Process collision counts
-positions = global_state_reshaped[:, :, 2:4]
-diffs = positions.unsqueeze(2) - positions.unsqueeze(1)
-pairwise_dist = torch.norm(diffs, dim=-1)
-collision_mask = (pairwise_dist <= 0.3) & ~torch.eye(n_agents, device=positions.device).bool().unsqueeze(0)
-collision_weights = collision_mask.sum(dim=-1).float()
-# Combine weights and softmax
-combined_weights = pos_weights - collision_weights
-softmax_weights = torch.softmax(combined_weights, dim=-1)
-global_q = (agents_q * softmax_weights).sum(dim=-1, keepdim=True)
-return global_q
-Compatibility Test.
-QLLM is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values.
-We used the TFCAF generated by QLLM to replace the mixing networks in three value-decomposition algorithms—RIIT, MASER, and QMIX.
-We then conducted experiments in the MPE environment to examine how integrating TFCAF affects the performance of these value-decomposition MARL methods.
-The experimental results tested in MPE environments are shown in Figure 5.
-Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
-Conclusion
+The primary LLM used in our method is the DeepSeek-R1 [7] inference model.
+For most experiments, the reported values represent the mean over five independent training runs, with 90% confidence intervals indicated by shaded areas.
+Notably, for SMAC scenarios, the solid lines represent the median test win rate, while the shaded regions denote the 25th-75th percentile interquartile range.
+Our experimental environments include: (1) Level-Based Foraging (LBF) [30], for testing in sparse reward settings; (2) Google Research Football (GRF) [16], to assess performance in complex, realistic environments; (3) Multi-agent Particle Environments (MPE) [23], for evaluating performance in high-dimensional state spaces; and (4) StarCraft Multi-Agent Challenge (SMAC) [36], to test performance in complex and highly cooperative multi-agent scenarios.
+We compare QLLM with seven baseline algorithms: QMIX [35], QPLEX [40], Qatten [43], RIIT [9], COMA [6], Weighted QMIX [34] and MASER [11].
+A detailed introduction and implementation details for all baselines and environments are provided in Appendix 0.C.
+5.2 The Superiority of QLLM
+Figure 3 show the performance of various algorithms in LBF and GRF environments. It is evident that the QLLM algorithm outperforms the baseline in both the LBF environment and the GRF environment.
+Figure 3: Performance comparison in LBF and GRF environments.
+5.3 The Generality of QLLM
+QLLM is based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-value.
+To verify this compatibility, we extend our evaluation beyond QMIX, using the TFCAF generated by QLLM to replace the mixing networks in two other value decomposition algorithms: RIIT and MASER.
+We conducted experiments in the MPE environment, and the experimental results are shown in Figure 4.
+Our results demonstrate that QLLM can be integrated with various value decomposition algorithms to significantly enhance their performance.
+Figure 4: Comparison of QLLM-enhanced variants with their original baselines in the MPE environments.
+5.4 The Performance of QLLM in High Dimensional State Spaces
+We evaluated the scalability of QLLM by increasing the number of agents to $N=15$ and $N=25$ in the simple-spread task.
+As shown in Figure 5, as state dimensionality grows, credit assignment becomes difficult for traditional value decomposition baselines, leading to significant performance degradation.
+In contrast, QLLM maintains high contribution assignment accuracy.
+This robustness stems from the LLM-generated nonlinear TFCAFs, which leverage well-defined task logic that is invariant to dimensionality, unlike traditional mixing network that struggle with high-dimensional parameter optimization.
+Figure 5: The results of high-dimensional testing in simple-spread task.
+5.5 The Performance on SMAC
+We evaluated QLLM on two hard SMAC maps: 3s_vs_5z and 2c_vs_64zg.
+As illustrated in Figure 6, QLLM outperforms all baselines with significantly faster convergence.
+These results demonstrate that LLM-generated logic efficiently solves complex coordination tasks where traditional neural mixers typically struggle to optimize during the early stages of training.
+Figure 6: Performance comparison on SMAC maps.
+5.6 The Interpretability of TFCAF
+To demonstrate the interpretability of QLLM, Listing 5.6 presents the complete generated TFCAF for the academy_3_vs_1_with_keeper scenario, while Figure 7 provides a visualization and a detailed explanation of its weight components.
+The generated logic explicitly leverages physical factors, such as ball possession and goal proximity, to attribute credit.
+Unlike traditional neural mixing networks that operate as black boxes, TFCAF offers a transparent and human-readable mechanism, clearly revealing the underlying tactical rationale for credit assignment in complex environments.
+{listing}
+[htbp] TFCAF for the academy_3_vs_1_with_keeper scenario in GRF.
+def QLLMNetwork(agents_q, global_state): # Extract environment informations ball_x, ball_y = global_state[:, 88], global_state[:, 89] possession, holder_my_team=global_state[:,94:97],global_state[:,97:101] ball_goal_dist = torch.sqrt((ball_x-1.0)**2 + ball_y**2) my_team_has_ball = torch.argmax(possession, dim=1) == 0 in_push_region = (ball_goal_dist>0.19) & (ball_goal_dist<0.99) & \ my_team_has_ball agent_x, agent_y = global_state[:,[0,2,4,6]],global_state[:,[1,3,5,7]] agent_goal_dist = torch.sqrt((agent_x-1.0)**2 + agent_y**2) agent_ball_dist = torch.sqrt((agent_x-ball_x.unsqueeze(1))**2 + (agent_y-ball_y.unsqueeze(1))**2) # Calculate weights components holder_weights = (1.0-ball_goal_dist).unsqueeze(1)*holder_my_team * \ in_push_region.unsqueeze(1)*10.0 non_push_weights = holder_my_team*my_team_has_ball.unsqueeze(1) * \ (~in_push_region).unsqueeze(1) support_weights=(1.0/(1.0+agent_goal_dist))*(~holder_my_team.bool())* \ my_team_has_ball.unsqueeze(1) defense_weights = (1.0/(1.0+agent_ball_dist)) * \ (~my_team_has_ball).unsqueeze(1) weights = torch.softmax(holder_weights + non_push_weights + support_weights + defense_weights, dim=1) return (agents_q * weights).sum(dim=1, keepdim=True)
+Figure 7: Visualization and Tactical Explanation of Weight Components in the TFCAF for academy_3_vs_1_with_keeper scenario.
+6 Conclusion
 In this study, we propose QLLM, a method that uses LLMs to address the credit assignment problem in multi-agent reinforcement learning through coder-evaluator framework.
-This approach generates training-free, task-specific credit assignment functions, reducing hallucination and improving task understanding.
-Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-based algorithms.
+This approach generates task-specific, training-free credit assignment functions, reducing hallucination and improving task understanding.
+Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-network-based algorithms.
 Future work will explore broader applications and deployment in real-world multi-robot systems.
-Appendix A LLM Prompts and Responses
-Below are the prompt template, example task information, and LLM response in our work.
-Role Prompt Coder’s ROLE Prompt: You are proficient in understanding tasks and writing Python code.
+Appendix 0.A LLM Prompts and Responses
+Coder Prompt You are proficient in understanding tasks and writing Python code.
 You should fully understand the multi-agent reinforcement learning tasks and environment, including the agents, action space, observation space, and global state space provided by the user.
@@
 Both input and output should be torch tensors.Individual Q-values represent each agent’s expected contribution to the team’s future rewards, while the global Q-value denotes the expected overall reward of the team as a whole.
-The input Q-values will have a shape of torch.Size([batchsize, n_agent]), the input global state will have a shape of torch.Size([batchsize, state_dim]), and the output global Q-value should have a shape of torch.Size([batchsize, 1]).You must strictly follow the input and output dimension requirements!
+The input Q-values will have a shape of torch.Size([batchsize, n_agent]), the input global state will have a shape of torch.Size([batchsize, state_dim]), and the output global Q-value should have a shape of torch.Size([batchsize, 1]).
+You must strictly follow the input and output dimension requirements!
 In the credit assignment function you wrote, You need to extract the individual components of the local observation vectors of the individual agents.
@@
 return global_q
-Evaluator’s ROLE Prompt: You are an excellent evaluator of multi-agent reinforcement learning performance.
+Evaluator Prompt You are an excellent evaluator of multi-agent reinforcement learning performance.
 I will attempt different functions in multi-agent reinforcement learning to solve the credit assignment problem.
@@
 Do not use any information that has not been provided to you to answer!
-Task Prompt Task Prompt of pz-mpe-simple-spread-v3: 1.
-Composition of the reward: We are playing Cooperative Navigation.
-In this task, 3 agents are asked to reach 3 landmarks.
-Each agent should move to get close to one landmark and avoid collision(distance<<=0.3) with other agents.
-The agent-team are rewarded based on how far the closest agent is to each landmark (the reward equals the negative value of the sum of the minimum distances).
-At the same time, the agent-team are penalized if they collide each other (-1 for each collision).
-At each step, each agent receives an observation array.
-2.
-State space: This observation is represented by an array with 18 dimensions: concat([agent’s velocity, agent’s location, 3 landmarks’ relative locations, 2 other agents’ relative locations, communication code]).
-The agent’s velocity is 2-dimensional, including x and y components.
-The agent’s location is 2-dimensional, including x and y coordinates.
-The landmarks’ relative locations are 3x2 dimensions, including relative x and y coordinates (landmark_i.x-agent.x, landmark_i.y-agent.y).
-The other agents’ relative locations are 2x2 dimensions, including relative x and y coordinates (agent_i.x-agent.x, agent_i.y-agent.y).
-The communication code are 4 dimensions, they have no role in credit assignment.
-The global state space is concatenate together from the observation space of the three agents, which are 18x3 dimensions.
-3.
-Action space: [move_left, move_right, move_down, move_up] Task Prompt of lbforaging:Foraging-8x8-2p-2f-coop-v3: The Level-Based Foraging environment focus on the coordination of involved agents.
+Task Prompt of LBF The Level-Based Foraging environment focus on the coordination of involved agents.
 The task for each agent is to navigate the grid-world map and collect items.
@@
 In order to collect a food item, agents have to choose a pick-up action next to the food item.
-However, such collection is only successful if 2 agents participate in the pickup at the same time.
+1.
+Composition of the reward: 8x8-2p-2f: Such collection is only successful if 2 agents participate in the pickup at the same time.
+The task is an 8x8 grid-world with 2 agents and 2 food items.
+10x10-3p-3f: Such collection is successful if the sum of involved agents’ levels is greater than or equal to the item level.
+The task is a 10x10 grid-world with 3 agents and 3 food items.
+15x15-4p-3f: Such collection is successful if the sum of involved agents’ levels is greater than or equal to the item level.
+The task is a 15x15 grid-world with 4 agents and 3 food items.
 At the same time the agents involved in the pickup must be one square above, below, left and right of the food item.
 Agents receive team reward equal to the level of the collected food item.
-In each step, every agent can choose to act by moving in one of four directions or attempt to pick up a food item.
-The task is an 8*8 grid-world with 2 agents and 2 food items scattered in the grid.
-Below, we will provide additional details to global states, actions and rewards for the Level-Based Foraging environment.
+Agents only receive non-zero team reward for picking up food items.
+2.
+States: Global state vector consists of 3 components: agents component, food items component and last step action.
+Total dimension: 8x8-2p-2f: 14 dimensions ($2\times 3+2\times 3+2$).
+10x10-3p-3f: 21 dimensions ($3\times 3+3\times 3+3$).
+15x15-4p-3f: 25 dimensions ($4\times 3+3\times 3+4$).
+Components detail: (1) Agents Component 8x8-2p-2f: 2 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level).
+10x10-3p-3f: 3 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level).
+15x15-4p-3f: 4 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level).
+(2) Food Items Component 8x8-2p-2f: 2 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level).
+10x10-3p-3f: 3 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level).
+15x15-4p-3f: 3 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level).
+When a food item has been picked up, then the respective values are replaced with (-1, -1, 0).
+(3) Last Actions Executed 8x8-2p-2f: 2 dims (This information is set to [-1, -1] when the environment is in the first step).
+10x10-3p-3f: 3 dims (This information is set to [-1, -1, -1] when the environment is in the first step).
+15x15-4p-3f: 4 dims (This information is set to [-1, -1, -1, -1] when the environment is in the first step).
+3.
+Actions: Each agent has six possible discrete actions: {0: Noop, 1: Move North, 2: Move South, 3: Move West, 4: Move East, 5: Pickup} While the first action corresponds to the action simply staying within its grid and the last action being used to pickup nearby food, the remaining four actions encode discrete 2D navigation.
+Picking up is only valid if the sum of the levels of the agents doing the action is greater than or equal to the level of the food.
+agents must be positioned one cell adjacent to the food.
+Task Prompt of GRF These tasks are offensive soccer games where the objective is to score goals by cooperating through passing and positioning to bring the ball as close as possible to the opponent’s goal, located at coordinates (1.0, 0.0).
 1.
-State space: Global state vector consists of 3 component(agents component,food items component and last step action)(Total dimension: 2x3+2x3+2=14 dims) 1.Agents Component (2 agents × 3 features = 6 dims) Feature 0: agent_i’s x coordinate (i=0,1) Feature 1: agent_i’s y coordinate Feature 2: agent_i’s level 2.Food Items Component (2 Food Items × 3 features = 6 dims) Feature 0: food_i’s x coordinate (i=0,1) Feature 1: food_i’s y coordinate Feature 2: food_i’s level When a food item has been picked up, then the respective values are replaced with (-1, -1, 0).You have to be careful to handle this particular situation properly.
-3.The actions just executed by the agents.
-(2 dims) {0,1,2…5} respectively represent {Noop, Move North, Move South, Move West, Move East, Pickup} This information is set to [-1 -1] when the environment is in the first step of each episode.
+Composition of the reward: The agents receive rewards consisting of two components: (1) Progress reward: Provides a team reward when an agent carries the ball closer to the opponent’s goal (granted if ball possession is maintained and ball-to-goal distance is between 0.19 and 0.99).
+(2) Goal reward: A team reward given when a goal is successfully scored, with an additional bonus based on the shot distance (farther distance yields a higher bonus).
+3_vs_1: Our team controls 1 goalkeeper and 3 offensive players (4 agents).
+The opposing team has 1 goalkeeper and 1 defender.
+Counterattack: This is an 11-vs-11 game.
+We control 4 players (4 agents), while the rest of our team and all opponents are controlled by built-in AI.
+Pass_and_Shoot: Our team consists of 1 goalkeeper and 2 offensive players (3 agents).
+The opposing team has 1 goalkeeper and 1 defender.
 2.
-Action space: In Level-Based Foraging environments, each agent has six possible discrete actions to choose at each timestep: Ai = {Noop, Move North, Move South, Move West, Move East, Pickup} While the first action corresponds to the action simply staying within its grid and the last action being used to pickup nearby food, the remaining four actions encode discrete 2D navigation.
-Upon choosing these actions, the agent moves a single cell in the chosen direction within the grid.
-Picking up is only valid if the sum of the levels of the agents doing the picking up action is greater than or equal to the level of the food.
-At the same time, the agents involved in the pickup must be positioned one cell above, below, to the left, or to the right of the food.
+States: Each reinforcement learning agent receives a local observation of 115 dimensions.
+The global state vector is formed by concatenating the local observations of the controlled agents.
+Components detail: (1) Players’ Positions and Directions 3_vs_1: 0–7: Positions of our 4 players; 8–15: Directions of our 4 players; 16–19: Positions of 2 opponent players; 20–23: Directions of 2 opponent players.
+Counterattack: 0–21: Positions of all our 11 players; 22–43: Directions of all our 11 players; 44–65: Positions of all 11 opponent players; 66–87: Directions of all 11 opponent players.
+Pass_and_Shoot: 0–5: Positions of our 3 players; 6–11: Directions of our 3 players; 12–15: Positions of 2 opponent players; 16–19: Directions of 2 opponent players.
+(2) Ball and Possession Information 88–90: Ball position (x, y, z); 91–93: Ball direction (x, y, z); 94–96: Ball possession (our team: [0, 1, 0], opponent team: [0, 0, 1], no one: [1, 0, 0]).
+(3) Agent Identification and Auxiliary Information 3_vs_1: 97–114: Invalid information.
+Counterattack: 97–107: 11-dimensional one-hot encoding indicating the agent’s ID (used to retrieve corresponding player’s data).
+108–114: Invalid information.
+Pass_and_Shoot: 97–114: Invalid information.
+Total Global State Dimension: The global state vector is constructed by concatenating the local observations of the three controlled agents, resulting in a total of 3_vs_1: 460 dimensions ($115\times 4$).
+Counterattack: 460 dimensions ($115\times 4$).
+Pass_and_Shoot: 345 dimensions ($115\times 3$).
+Task Prompt of MPE 1.
+Composition of the reward: Spread: We are playing Cooperative Navigation.
+In this task, NN agents are asked to reach NN landmarks (where $N\in\{6,15,25\}$).
+Each agent should move to get close to one landmark and avoid collision (distance $\leq$ 0.3) with other agents.
+The agent-team are rewarded based on how far the closest agent is to each landmark (the reward equals the negative value of the sum of the minimum distances).
+At the same time, the agent-team are penalized if they collide each other (-1 for each collision).
+Adversary: In this environment, there is 1 adversary, 2 good agents, 2 landmarks.
+All agents observe the position of 2 landmarks and other agents.
+One of these 2 landmarks is the ’target landmark’.
+Good agents are rewarded based on how close the closest one of them is to the target landmark, but negatively rewarded based on how close the adversary is to the target landmark.
+The adversary want to get closer to the target landmark based on the agents’ behavior, but it doesn’t know which landmark is the target landmark.
+Tag: We are playing Predator Prey.
+In this task, 3 agents should cooperate to catch 1 prey.
+Agents move to get closer to the prey to catch it (distance << 0.125).
+Prey is faster and tries to get out of the agents’ clutches.
+Agents are slower and are rewarded for hitting preys (+10 for each collision).
+Obstacles will block the way.
+At each step, each agent receives an observation array.
+2.
+State space: This observation is represented by an array with: Spread: 36 dimensions (for $N=6$), 90 dimensions (for $N=15$), or 150 dimensions (for $N=25$).
+Adversary: 8 dimensions.
+Tag: 16 dimensions.
+The observation components are concatenated as: Spread: concat([agent’s velocity, agent’s location, NN landmarks’ relative locations, N1N-1 other agents’ relative locations, communication code]).
+Adversary: concat([target landmark’s relative locations, 2 landmarks’ relative locations, other agents’ relative locations (including adversary)]).
+Tag: concat([agent’s velocity, agent’s location, 2 obstacles’ relative locations, 2 other agents’ relative locations, 1 prey’s relative locations, 1 prey’s velocity]).
+The global state space is concatenated together from the observation space of all agents, resulting in: Spread: $36\times 6$ (for $N=6$), $90\times 15$ (for $N=15$), or $150\times 25$ (for $N=25$) dimensions.
+Adversary: $8\times 2$ dimensions.
+Tag: $16\times 3$ dimensions.
 3.
-Composition of the reward: Agents only receive non-zero team reward for picking up food items within the Level-Based Foraging environment.
-The reward for picking up a food item equals to the food level.
-Task Prompt of matrixgames:climbing-nostate-v0: This task is a two-player cooperative matrix games with a given 3×3 matrix: [[11, -30, 0], [-30, 7, 0], [0, 6, 5]] 1.
-Action space: The environment consists of two agents, agent_0 and agent_1, each having three possible actions(0,1,2) corresponding to the row and column indices of the matrix.
-2.
-Composition of the reward: Once both agents select their actions, they receive a shared reward determined by the corresponding entry in the matrix.
-There is no communication between the two agents, and the objective is to maximize the shared reward.
-Be sure to prevent the strategy from falling into a local optimum, and only get the highest reward if both intelligences choose action 0.
-3.
-State space: The state vector is 2 dims, storing the actions that the two agents have just executed.
-At the first step at the beginning of an episode, the state vector is [-1,-1], because the agents have not yet started executing actions.
-Please take care to handle this special case.
-Task Prompt of academy_counterattack_easy: 1.
-Action space: This is an offensive football game where our team controls four reinforcement learning agents, including one goalkeeper and three attacking players.
-The opposing team consists of one goalkeeper and one defender.
-The objective is to move the ball as close as possible to the opponent’s goal, located at coordinate (1.0, 0.0), through passing and off-ball movement, and to score a goal.
-2.
-Composition of the reward: The agents receive rewards based on two components: progress reward and goal-scoring reward.
-The progress reward is given when a player on our team carries the ball closer to the opponent’s goal, providing a team reward.
-This reward is only available when our player is in possession of the ball and the ball’s distance to the opponent’s goal is greater than 0.19 and less than 0.99.
-The player in possession is usually the one closest to the ball.
-The goal-scoring reward is provided when our team successfully scores.
-In addition to this reward, an extra reward is given based on the distance from which the goal is scored.
-The farther the distance, the higher the extra reward.
-3.
-State space: Each agent has a local observation vector of 115 dimensions: 0–7: positions of our four players, each represented by (x, y) 8–15: directions of our four players, each represented by (x, y) 16–19: positions of two opponent players, each represented by (x, y) 20–23: directions of two opponent players, each represented by (x, y) 24–87: invalid information 88–90: ball position (x, y, z) 91–93: ball direction (x, y, z) 94–96: ball possession, encoded as one-hot vectors: our team [0, 1, 0], opponent [0, 0, 1], no possession [1, 0, 0] 97–114: invalid information The global state vector is constructed by concatenating the local observations of all four agents, resulting in a total of 460 dimensions.
+Action space: [no_action, move_left, move_right, move_down, move_up]
+Task Prompt of SMAC SECTION 1: CONTEXT AND OBJECTIVE You are an expert in Multi-Agent Reinforcement Learning (MARL), specifically focusing on the QLLM framework.
+Your task is to generate a PyTorch function named QLLMNetwork(agents_q, state).
+This function must act as a non-linear mixing network that takes individual agent Q-values with shapes: 3s_vs_5z: agents_q, shape [Batch, 3] 2c_vs_64zg: agents_q, shape [Batch, 2] and the global state with shapes: 3s_vs_5z: state, shape [Batch, 68] 2c_vs_64zg: state, shape [Batch, 342] to compute a scalar Global Q-value (shape [Batch, 1]).
+You must utilize the detailed features in the state vector to dynamically compute weights and biases for summing the agent Q-values.
+The agents receive rewards consisting of damage rewards (+1 per damage), kill rewards (+10 per kill), and a win bonus (+200).
+The return is scaled such that the maximum possible scaled return is 20.
+SECTION 2: ENVIRONMENT SPECIFICATION 3s_vs_5z: 3 Stalkers vs 5 Zealots.
+Stalkers (Allies) have 80 HP, 80 Shield, and 0.2142 attack range.
+Zealots (Enemies) have 100 HP, 50 Shield, and melee range (0.05).
+2c_vs_64zg: 2 Colossi vs 64 Zerglings.
+Colossi (Allies) have 200 HP, 150 Shield, and 0.25 attack range with splash damage.
+Zerglings (Enemies) have 35 HP, 0 Shield, and melee range (0.01).
+Coordinates (X, Y) are normalized relative to the map center (range [-0.5, 0.5]) using a coordinate formula: (absolute_coord - map_center) / 28.
+SECTION 2.5: HEURISTIC REASONING GUIDE 3s_vs_5z: Stalkers are ranged units facing melee Zealots.
+Success depends on “kiting"—maintaining a distance between 0.05 and 0.2142.
+Agents successfully kiting while attacking should receive higher weights.
+2c_vs_64zg: Colossi deal powerful splash damage.
+Success depends on maintaining distance from the massive swarm of 64 Zerglings.
+Credit should be weighted toward agents effectively engaging clusters of enemies while kiting.
+An agent being swarmed (multiple enemies at melee distance) should have its survival actions prioritized.
+Common Rules: (1) Mask Dead Agent: If health is 0.0, weight must be forced to $0.$ (2) Efficiency: Agents with high weapon cooldown cannot attack; temporarily lower their weights unless they are drawing fire.
+(3) Shield Awareness: Protoss survivability depends on “Health + Shield".
+Priority should be given to units with low total survivability.
+SECTION 3: DATA STRUCTURE PART A: ALLY INFORMATION 3s_vs_5z: Indices 0-14.
+3 agents, 5 features each: [Health, Cooldown, Rel_X, Rel_Y, Shield].
+2c_vs_64zg: Indices 0-9.
+2 agents, 5 features each: [Health, Cooldown, Rel_X, Rel_Y, Shield].
+PART B: ENEMY INFORMATION 3s_vs_5z: Indices 15-34.
+5 enemies, 4 features each: [Health, Rel_X, Rel_Y, Shield].
+2c_vs_64zg: Indices 10-201.
+64 enemies, 3 features each: [Health, Rel_X, Rel_Y].
+PART C: LAST ACTIONS 3s_vs_5z: Indices 35-67.
+3 agents, 11 actions each.
+Actions: 0:No-op, 1:Stop, 2-5:Move, 6-10:Attack Enemy 1-5.
+2c_vs_64zg: Indices 202-341.
+2 agents, 70 actions each.
+Actions: 0:No-op, 1:Stop, 2-5:Move, 6-69:Attack Enemy 1-64.
+Please write the QLLMNetwork function in PyTorch, ensuring all tensors are on state.device.
 TFCAF template import torch import torch.nn as $n$ import torch.nn.functional as F import numpy as np class QLLMMixer($n$.Module): def __init__(self): super(QLLMMixer, self).__init__() def forward(self, agents_q, global_state): agents_q = torch.squeeze(agents_q) a = agents_q.shape[0] b = agents_q.shape[1] agents_q = agents_q.reshape(-1, agents_q.shape[-1]) global_state = global_state.reshape(-1, global_state.shape[-1]) #This is generated by LLM return (global_qagents_q.shape[-1]).reshape(a, b, 1).cuda()
-LLM Generated Responses LLM Response of pz-mpe-simple-spread: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batchsize = agents_q.size(0) n_agents = 6 global_state_reshaped = global_state.view(batchsize, n_agents, 36) # Process landmark distances landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2) landmark_distances = torch.norm(landmarks_relative, dim=-1) min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1) pos_weights = 1.0 / min_dist_per_agent # Process collision counts positions = global_state_reshaped[:, :, 2:4] diffs = positions.unsqueeze(2) - positions.unsqueeze(1) pairwise_dist = torch.norm(diffs, dim=-1) collision_mask = (pairwise_dist <<= 0.3)  &  torch.eye(n_agents, device=positions.device).bool().unsqueeze(0) collision_weights = collision_mask.sum(dim=-1).float() # Combine weights and softmax combined_weights = pos_weights - collision_weights softmax_weights = torch.softmax(combined_weights, dim=-1) global_q = (agents_q  softmax_weights).sum(dim=-1, keepdim=True) return global_q LLM Response of lbforaging:Foraging-8x8-2p-2f-coop-v3: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) agents_data = global_state[:, :6].view(batch_size, 2, 3) food_data = global_state[:, 6:12].view(batch_size, 2, 3) agents_level = agents_data[:, :, 2] agents_pos = agents_data[:, :, :2].unsqueeze(2) food_pos = food_data[:, :, :2].unsqueeze(1) dxdy = torch.abs(agents_pos - food_pos).sum(dim=-1) adjacent = (dxdy == 1).float() active_food_mask = ((food_data[:, :, 0] != -1) & (food_data[:, :, 1] != -1)).unsqueeze(1) adjacent = adjacent  active_food_mask total_contribution = torch.zeros_like(agents_level) for j in range(2): adjacent_j = adjacent[:, :, j] food_level_j = food_data[:, j, 2] sum_adj_levels = (agents_level  adjacent_j).sum(dim=1) num_adj = adjacent_j.sum(dim=1) valid = ((num_adj >>= 2) & (sum_adj_levels >>= food_level_j)).float().unsqueeze(1) contribution_j = adjacent_j  food_level_j.unsqueeze(1)  valid total_contribution += contribution_j agent_contrib = agents_level  total_contribution weights = torch.softmax(agent_contrib, dim=1) global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of academy_counterattack_easy: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) n_agents = agents_q.size(1) # Extract ball information from first agent’s observation ball_pos = global_state[:, 88:90] holder_info = global_state[:, 94:97] is_my_team = holder_info[:, 1].bool() # Calculate distance to goal goal_pos = torch.tensor([1.0, 0.0], device=agents_q.device) ball_goal_dist = torch.norm(ball_pos - goal_pos, dim=1) valid_advance = is_my_team & (ball_goal_dist >> 0.19) & (ball_goal_dist << 0.99) # Calculate agent-ball distances and holder mask all_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-ball_pos[:,0], pos_y-ball_pos[:,1]], dim=1), dim=1) all_dists.append(dist.unsqueeze(1)) dist_matrix = torch.cat(all_dists, dim=1) holder_mask = F.one_hot(torch.argmin(dist_matrix, dim=1), num_classes=n_agents).bool() # Calculate agent-goal distances goal_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-goal_pos[0], pos_y-goal_pos[1]], dim=1), dim=1) goal_dists.append(dist.unsqueeze(1)) goal_dist_matrix = torch.cat(goal_dists, dim=1) # Calculate weights components advance_component = holder_mask.float()  valid_advance.unsqueeze(1)  5.0 position_component = 1.0 / (goal_dist_matrix + 1e-6) # Combine components with softmax combined = torch.where(valid_advance.unsqueeze(1), advance_component, position_component) weights = F.softmax(combined, dim=1) # Compute global Q-value global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of matrixgames:climbing-nostate-v0: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: device = agents_q.device batch_size = agents_q.shape[0] # Initialize base weights and reward matrix weights = torch.full((batch_size, 2), 0.5, device=device) reward_matrix = torch.tensor([[11, -30, 0], [-30, 7, 0], [0, 6, 5]], device=device, dtype=torch.float32) # Create mask for valid states (non-initial) valid_mask = (global_state[:, 0] >>= 0) & (global_state[:, 1] >>= 0) if valid_mask.any(): prev_actions = global_state[valid_mask].long() a0 = prev_actions[:, 0] a1 = prev_actions[:, 1] # Calculate previous rewards and optimal condition prev_rewards = reward_matrix[a0, a1] optimal_mask = (prev_rewards == 11) # Calculate deviation from optimal for each agent agent0_deviation = (a0 != 0).float() agent1_deviation = (a1 != 0).float() # Create penalty weights based on deviation penalty_weights = torch.stack([1.0 - agent0_deviation  0.5,1.0 - agent1_deviation  0.5], dim=1) # Apply optimal condition override penalty_weights[optimal_mask] = 0.5 # Maintain equal weights for optimal case # Normalize weights to maintain sum=1 weight_sums = penalty_weights.sum(dim=1, keepdim=True) normalized_weights = penalty_weights / weight_sums # Update weights for valid samples weights[valid_mask] = normalized_weights # Compute global Q-value with dynamic weighting global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q
-Appendix B Details of Environments and Baselines
-Baseline Algorithms
+Appendix 0.B Additional Experiments
+0.B.1 Ablation Study on the Evaluator LLM
+While most LLM-enhanced reinforcement learning methods rely on a single model, our coder-evaluator framework leverages the collaboration between $M_{\text{coder}}$ and $M_{\text{evaluator}}$ to ensure logical precision.
+To isolate the impact of the evaluator, we evaluated a variant termed QLLM-C, which utilizes only the coder LLM for single-shot generation (effectively $K=1,T=1$).
+We conducted this ablation study on the SMAC 3s_vs_5z and 2c_vs_64zg scenarios.
+As illustrated in Figure 8, QLLM-C exhibits a slight performance advantage over the QMIX baseline, suggesting that the zero-shot prior knowledge encoded in LLMs can capture basic coordination patterns.
+However, QLLM-C is significantly outperformed by the full QLLM framework in both scenarios.
+This performance gap demonstrates that $M_{\text{evaluator}}$ is essential for filtering out hallucinations and refining credit assignment logic through iterative feedback, especially in complex environments where high-fidelity reasoning is critical.
+These results validate the architectural necessity of our coder-evaluator framework.
+Figure 8: Performance comparison of QLLM and QLLM-C on SMAC maps. QLLM-C stands for the variant utilizing only the coder LLM for generation.
+0.B.2 Ablation Study on Candidate Number and Iteration Rounds
+To investigate the impact of the number of candidate TFCAFs per round ($K$) and the number of iteration rounds within the coder-evaluator framework ($T$), we conducted parameter ablation studies on the MPE simple-tag task.
+As illustrated in Figure 9, the results reveal that in relatively simple scenarios like MPE, generating a single TFCAF without iterative optimization (i.e., $K=1,T=1$) is already sufficient to achieve near-optimal performance.
+However, this finding contrasts with our observations in complex environments such as SMAC.
+In fact, the QLLM-C variant (refer to Fig.
+8) effectively represents the $K=1,T=1$ configuration, and its performance in SMAC scenarios is significantly inferior to that of the standard $QLLM(K=3,T=3)$.
+We can conclude that for low-complexity tasks, minimal generation counts and iterations are adequate.
+In contrast, complex scenarios necessitate a larger candidate pool and multiple refinement cycles to synthesize TFCAFs with sophisticated and robust coordination logic.
+Figure 9: Ablation study of the number of candidates $K$ and iteration rounds $T$ on the pz-mpe-simple-tag scenario.
+0.B.3 Compatibility Across Different LLMs
+To assess framework compatibility, we implemented QLLM using DeepSeek-R1 [7], DeepSeek-V3111https://www.deepseek.com/, and Mistral-Large [^2] https://mistral.ai/ on the MPE simple-tag task.
+As illustrated in Fig.
+10, QLLM consistently surpasses the QMIX baseline regardless of the underlying LLM architecture.
+Furthermore, the results show a positive correlation between LLM reasoning capabilities and RL performance; specifically, the Mistral-Large variant exhibits a slight performance margin over the others.
+These findings demonstrate that QLLM is a versatile framework capable of effectively leveraging the diverse reasoning strengths of various state-of-the-art LLMs.
+Figure 10: Performance comparison of QLLM variants using different mainstream LLMs on the MPE simple-tag scenario.
+0.B.4 Comparison of Algorithms Training Costs
+As discussed in the methodology section, QLLM replaces the traditional neural mixing network with a deterministic functional expression, TFCAF.
+This architectural shift results in a significantly lower number of learnable parameters compared to standard value decomposition methods.
+Table 1 details the specific number of trainable parameters for each algorithm across the tested environments, while Figure 11 provides a visual comparison of the average parameter counts.
+As illustrated, QLLM reduces the number of trainable parameters by approximately 13% to 37% relative to the baseline average.
+This substantial reduction simplifies the model structure and directly lowers the training cost, thereby enhancing the overall efficiency of the algorithm while maintaining superior performance.
+Furthermore, we evaluated the operational efficiency of QLLM by comparing its average training duration against QMIX across these SMAC scenarios.
+Experimental results show that QLLM reduces the average training time per step from 161.00 ms to 104.18 ms (35.3% reduction) and shortens the average total duration for 2 million training steps from approximately 9.05 hours to 5.38 hours (40.5% reduction).
+Although QLLM includes an additional phase for TFCAF synthesis, one iteration of the code generation process (with $n=3$) takes only about 15 minutes, which is negligible compared to the several hours of training time saved.
+These findings demonstrate that QLLM not only achieves superior performance but also significantly optimizes training costs and time efficiency by reducing learnable parameters and accelerating the training procedure.
+Table 1: Number of learnable 0(in K) for each environment under different algorithms
+Environment / Algorithm QLLM QMIX QPLEX Qatten RIIT Baseline Average Reduction Rate
+pz-mpe-simple-spread 105.22 140.42 154.76 191.11 191.18 166.74 36.88%
+pz-mpe-simple-tag 102.28 112.81 128.01 129.03 132.20 125.91 18.78%
+pz-mpe-simple-adversary 101.38 107.27 122.76 118.28 124.36 118.58 14.52%
+Foraging-2s-15x15-4p-3f 103.17 111.82 127.37 121.96 130.54 123.28 16.30%
+Foraging-2s-8x8-2p-2f-coop 101.77 106.89 122.76 116.68 124.30 118.15 13.87%
+Foraging-2s-10x10-3p-3f 102.66 109.74 125.45 120.04 127.91 121.19 15.28%
+academy_3_vs_1_with_keeper 430.36 529.30 543.01 601.50 568.09 575.36 25.23%
+academy_counterattack_easy 430.36 529.30 543.01 601.50 568.09 575.36 25.23%
+academy_pass_and_shoot_with_keeper 430.10 504.88 523.05 560.76 530.87 543.13 20.83%
+Figure 11: Comparison of the average number of learnable parameters across all tested environments.
+0.B.5 Experimental Setup
+The experiments were conducted on computing nodes equipped with an Intel(R) Xeon(R) Platinum 8488C CPU and NVIDIA H100 PCIe GPUs with 80GB of VRAM.
+Our framework was implemented using Python 3.8.20, with PyTorch 2.1.0 serving as the deep learning backend.
+To ensure high-performance GPU acceleration, the system was configured with NVIDIA Driver version 570.195.03 and CUDA 12.8.
+Appendix 0.C Details of Environments and Baselines
+0.C.1 Baseline Algorithms
 QMIX (Rashid et al.
@@
 This design expands the class of joint value functions that can be represented under the CTDE paradigm.
-Weighted QMIX (Rashid et al.
-2020a) further generalizes QMIX by learning an additional weighting factor for each agent’s contribution to the global Q-value.
-By relaxing the monotonicity assumption through these learnable weights, it improves the accuracy of credit assignment in tasks with more complex cooperative dependencies.
 Qatten (Yang et al.
@@
 This counterfactual reasoning helps reduce variance in the policy gradient and promotes coordinated behaviors.
+Weighted QMIX (Rashid et al.
+2020a) further generalizes QMIX by learning an additional weighting factor for each agent’s contribution to the global Q-value.
+By relaxing the monotonicity assumption through these learnable weights, it improves the accuracy of credit assignment in tasks with more complex cooperative dependencies.
 MASER (Jeon et al.
@@
 By augmenting standard Q-learning with subgoal-conditioned value estimation, MASER enhances exploration efficiency and coordination while remaining compatible with centralized training and decentralized execution.
-Code Repository
-The baseline algorithm in this experiment is implemented based on the epymarl [^5] https://github.com/uoe-agents/epymarl and pymarl2666https://github.com/hijkzzz/pymarl2 codebases.
-We tested MPE, LBF, and matrix games using the epymarl library, while Google Research Football was evaluated with pymarl2.
-Although the SMAC [^7] https://github.com/oxwhirl/smac environment is a popular benchmark for multi-agent reinforcement learning and is built upon PySC2888https://github.com/google-deepmind/pysc2, it lacks comprehensive documentation regarding the specific unit types, abilities, and map configurations used in each scenario.
-This obscurity hinders reproducibility and limits fine-grained control over experimental variables.
-Consequently, we opted for several environments with better transparency and configurability to ensure experimental rigor and facilitate controlled evaluations.
-Tasks Details
-Matrix Games
-The cooperative matrix games environment is a simple benchmark setting in multi-agent reinforcement learning, where two or more agents simultaneously select actions from predefined action spaces.
-The environment is defined by a payoff matrix that specifies the global reward received based on the joint actions of all agents (below is the payoff matrix used in test environment).
-Cooperation is essential, as the optimal strategy often depends on coordinated decision-making among agents.
-In order to make sense of the global state space of this task, we include the actions executed by the agents last step.
-$\begin{bmatrix}0&6&5\\ -30&7&0\\ 11&-30&0\end{bmatrix}$
-Figure 7: Environment visualizations of Level-Based Foraging (LBF), Google Research Football (GRF), and Multi-Agent Particle Environments (MPE)
-Level-Based Foraging
+0.C.2 Code Repository
+The baseline algorithms in this experiment are implemented based on the epymarl [^3] https://github.com/uoe-agents/epymarl, pymarl [^4] https://github.com/oxwhirl/pymarl, and pymarl2555https://github.com/hijkzzz/pymarl2 codebases.
+Specifically, we evaluated the MPE and LBF environments using the epymarl library.
+The StarCraft Multi-Agent Challenge (SMAC) was tested utilizing the original pymarl framework, while Google Research Football (GRF) was evaluated with pymarl2.
+This setup ensures experimental rigor and facilitates fair and controlled comparisons across diverse cooperative tasks.
+0.C.3 Tasks Details
+Figure 12: Environment visualizations of Level-Based Foraging (LBF), Google Research Football (GRF), Multi-Agent Particle Environments (MPE) and StarCraft Multi-Agent Challenge (SMAC).
+0.C.3.1 Level-Based Foraging
 The Level-Based Foraging environment consists of tasks focusing on the coordination of involved agents.
@@
 Foraging-2s-10x10-3p-3f 3 21 6
-Multi-Agent Particle Environments
+0.C.3.2 Google Research Football
+The Google Research Football environment is a challenging reinforcement learning platform designed to simulate football matches in a highly realistic and strategic setting.
+Agents control individual players and must learn to cooperate, plan, and execute complex behaviors to achieve objectives such as scoring goals or defending against opponents.
+The environment provides various predefined scenarios as well as full 11-versus-11 matches, supporting both single-agent and multi-agent training.
+In the Google Research Football environment, the action space consists of a discrete set of predefined high-level actions such as moving, passing, shooting, and sprinting, allowing agents to focus on strategic decision-making rather than low-level motor control.
+The Simple115 observation space is a fixed-size vector with 115 features, providing essential information about the game state, including the positions and velocities of the ball and players, the directions of movement, and other key attributes necessary for learning effective policies.
+Table 3: Details of GRF tasks
+GRF Tasks N_agents State Shape Action Num
+academy_3_vs_1_with_keeper 4 460 19
+academy_counterattack_easy 4 460 19
+academy_pass_and_shoot_with_keeper 3 345 19
+0.C.3.3 Multi-Agent Particle Environments
+The Multi-Agent Particle Environments are a widely-used set of communication-oriented environments where particle agents can move, communicate, observe each other, exert physical forces, and interact with fixed landmarks within a 2D continuous space[1].
+These environments are highly configurable and serve as standard benchmarks for evaluating coordination, competition, and communication in MARL.
 Simple-Spread: In this task, three agents are trained to move to three landmarks while avoiding collisions with each other.
@@
 pz-mpe-simple-adversary 2 16 5
-Google Research Football
-The Google Research Football environment is a challenging reinforcement learning platform designed to simulate football matches in a highly realistic and strategic setting.
-Agents control individual players and must learn to cooperate, plan, and execute complex behaviors to achieve objectives such as scoring goals or defending against opponents.
-The environment provides various predefined scenarios as well as full 11-versus-11 matches, supporting both single-agent and multi-agent training.
-In the Google Research Football environment, the action space consists of a discrete set of predefined high-level actions such as moving, passing, shooting, and sprinting, allowing agents to focus on strategic decision-making rather than low-level motor control.
-The Simple115 observation space is a fixed-size vector with 115 features, providing essential information about the game state, including the positions and velocities of the ball and players, the directions of movement, and other key attributes necessary for learning effective policies.
-Table 3: Details of GRF tasks
-GRF tasks n_agents state shape action num
-academy_3_vs_1_with_keeper 4 460 19
-academy_counterattack_easy 4 460 19
-academy_pass_and_shoot_with_keeper 3 345 19
-Appendix C Additional Experiments
-Compatibility Testing of Different LLMs
-We tested the performance of QLLM with three mainstream LLMs, Deepseek-R1, Deepseek-V3999https://www.deepseek.com/, and ChatGPT-4o [^10] https://openai.com/, along with other algorithms in the MPE environment.
-As shown in Figure 8, the experimental results indicate that mainstream LLMs outperform the baseline algorithms.
-Moreover, as the inference capabilities of the LLMs improve, the performance of our algorithm also improves.
-Figure 8: Performance comparison of QLLM using different LLMs in the GRF environment.
-Tests of the Effect of the Evaluator LLM
-Most reinforcement learning algorithms enhanced by LLMs use only one kind of LLM to do their job directly.
-Our coder-evaluator framework features two LLMs, each with its specific role, working together to generate an accurate credit assignment function.
-Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in LLMs.
-We tested the performance of QLLM across different reinforcement learning environments without using the evaluator LLM.
-The experimental results (Figure 9) show that, the evaluator LLM significantly improves the algorithm’s performance in most environments.
-Figure 9: Performance comparison of QLLM without evaluator LLM in the GRF environment, where QLLM-C stands for QLLM that do not use the evaluator LLM
-Comparison of Algorithms Training Costs
-As mentioned in the main text, QLLM replaces the mixing network with a deterministic functional expression, TFCAF, resulting in a significantly smaller number of learnable parameters.
-Therefore, the training cost of our algorithm is reduced.
-Table 4 lists the number of trainable parameters for each algorithm in the environments tested in this paper.
-Our algorithm reduces the number of trainable parameters by approximately 12% to 37%.
-This reduction directly lowers the training cost and simplifies the model structure, making the algorithm more efficient and improving performance.
-Table 4: Number of learnable parameters (in K) for each environment under different algorithms
-Environment / Algorithm QLLM QMIX QATTEN MADDPG RIIT QPLEX Baseline Average Reduction Rate (%)
-pz-mpe-simple-spread 105.221 140.422 191.114 154.246 191.175 154.762 166.744 36.88%
-pz-mpe-simple-tag 102.277 112.806 129.034 127.494 132.199 128.010 125.909 18.78%
-pz-mpe-simple-adversary 101.381 107.270 118.282 122.246 124.359 122.762 118.584 14.52%
-Foraging-2s-15x15-4p-3f 103.174 111.815 121.963 126.727 130.536 127.372 123.283 16.30%
-Foraging-2s-8x8-2p-2f-coop 101.766 106.887 116.683 122.119 124.296 122.764 118.150 13.87%
-Foraging-2s-10x10-3p-3f 102.662 109.735 120.043 124.807 127.912 125.452 121.190 15.28%
-climbing-nostate-v0 100.099 103.684 110.408 118.148 118.661 118.406 113.861 12.07%
-academy_3_vs_1_with_keeper 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
-academy_counterattack_easy 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
-academy_pass_and_shoot_with_keeper 430.099 504.884 560.760 600.084 530.869 523.046 543.129 20.83%
+0.C.3.4 StarCraft Multi-Agent Challenge
+The StarCraft Multi-Agent Challenge is a high-dimensional micro-management benchmark based on the StarCraft II engine.
+In this environment, agents must learn to cooperate with each other and execute sophisticated combat strategies (such as focus fire and kiting) to defeat enemy units and achieve victory.
+To evaluate the performance of our algorithm in complex multi-agent cooperative scenarios, we selected two challenging maps categorized as hard difficulty: 3s_vs_5z and 2c_vs_64zg.
+These scenarios require sophisticated micromanagement and coordination to handle heterogeneous units or large-scale enemy swarms.
+The specific parameters of these two maps are detailed in Table 5.
+Table 5: Details of the SMAC scenarios.
+SMAC Tasks Ally Units Enemy Units State Shape Action Num
+3s_vs_5z 3 Stalkers 5 Zealots 68 11
+2c_vs_64zg 2 Colossi 64 Zerglings 342 70
+Appendix 0.D Proof of the QLLM Decomposition Formula
+In this section, we provide a formal derivation of the decomposition formulation used in QLLM (Equation (3)).
+We aim to prove that the joint action-value function $Q_{\text{tot}}(s,\boldsymbol{a})$ can be represented as a state-dependent linear combination of individual utility functions $Q_{i}(\tau^{i},a^{i})$ plus a state-dependent bias term $f_{b}(s)$.
+Assuming the joint value function $Q_{\text{tot}}$ is differentiable with respect to the individual $Q$-values, we can perform a multivariate Taylor expansion of $Q_{\text{tot}}$ at the state $s$ as a function of the utility vector $\boldsymbol{Q}=(Q_{1},\dots,Q_{n})$:
+$Q_{\text{tot}}(s,\boldsymbol{Q})=\text{const}(s)+\sum_{i=1}^{n}\mu_{i}(s)Q_{i}+\sum_{i,j}\mu_{ij}(s)Q_{i}Q_{j}+\mathcal{O}(Q^{3}),$ (10)
+where
+$\mu_{i}(s)=\frac{\partial Q_{\text{tot}}}{\partial Q_{i}}\quad\text{and}\quad\mu_{ij}(s)=\frac{1}{2}\frac{\partial^{2}Q_{\text{tot}}}{\partial Q_{i}\partial Q_{j}}.$
+To analyze the local behavior near an optimal joint action $\boldsymbol{a}_{0}=(a_{0}^{1},\dots,a_{0}^{n})$, we note that the gradient $\nabla_{\boldsymbol{a}}Q_{\text{tot}}$ vanishes at $\boldsymbol{a}_{0}$.
+Since
+$\frac{\partial Q_{\text{tot}}}{\partial a^{i}}=\frac{\partial Q_{\text{tot}}}{\partial Q_{i}}\frac{\partial Q_{i}}{\partial a^{i}}=0,$
+and assuming that each agent has a non-trivial impact on the team ($\mu_{i}\neq 0$), the local gradient of each agent’s utility must also vanish:
+$\frac{\partial Q_{i}}{\partial a^{i}}(a_{0}^{i})=0.$
+Consequently, each individual utility function can be approximated by its local quadratic expansion:
+$Q_{i}(a^{i})\approx\alpha_{i}(s)+\beta_{i}(s)(a^{i}-a_{0}^{i})^{2}.$
+The key to linearizing the formulation lies in the treatment of the second-order cross terms $Q_{i}Q_{j}$. Substituting the quadratic approximation into these terms yields:
+$Q_{i}Q_{j}$ $\approx[\alpha_{i}+\beta_{i}(a^{i}-a_{0}^{i})^{2}][\alpha_{j}+\beta_{j}(a^{j}-a_{0}^{j})^{2}]$
+$=\alpha_{i}\alpha_{j}+\alpha_{i}\beta_{j}(a^{j}-a_{0}^{j})^{2}+\alpha_{j}\beta_{i}(a^{i}-a_{0}^{i})^{2}+\textit{higher-order terms}$
+$\approx\alpha_{i}\alpha_{j}+\alpha_{i}(Q_{j}-\alpha_{j})+\alpha_{j}(Q_{i}-\alpha_{i})$
+$=\alpha_{j}Q_{i}+\alpha_{i}Q_{j}-\alpha_{i}\alpha_{j}.$ (11)
+Equation (11) demonstrates that the product of two utilities near the optimum can be effectively represented as a linear combination of $Q_{i}$ and $Q_{j}$ plus a state-dependent constant.
+By substituting this result into the Taylor expansion in Equation (10) and grouping all terms associated with each $Q_{i}$, we obtain:
+$\begin{split}Q_{\text{tot}}(s,\boldsymbol{a})&\approx\sum_{i=1}^{n}\underbrace{\left[\mu_{i}(s)+2\sum_{j\neq i}\mu_{ij}(s)\alpha_{j}(s)+\dots\right]}_{f_{w}^{i}(s)}Q_{i}\\ &\quad+\underbrace{\left[\text{const}(s)-\sum_{i,j}\mu_{ij}(s)\alpha_{i}\alpha_{j}+\dots\right]}_{f_{b}(s)}.\end{split}$ (12)
+This derivation confirms that the formulation $Q_{\text{tot}}(s,\boldsymbol{a})=\sum f_{w}^{i}(s)Q_{i}+f_{b}(s)$ is a mathematically sound approximation for any complex value-based mixing mechanism that satisfies the Individual-Global-Max (IGM) principle.
+While traditional methods like QMIX or Qatten utilize hypernetworks to estimate these coefficients, QLLM leverages the logical reasoning of LLMs to generate the functions $f_{w}^{i}(s)$ and $f_{b}(s)$ directly, which enhances both interpretability and parameter efficiency.